### v1

# PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud by 2D Inpainting

Abstract.
Reconstructing textured meshes from colored point clouds is an important but challenging task in 3D graphics and vision. Most existing methods predict colors as implicit functions in 3D or UV space, suffering from blurry textures or the lack of generalization capability. Addressing this, we propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud. It produces meshes with enhanced fidelity and clarity by 2D image inpainting, taking advantage of the mature techniques and massive data of 2D vision. Specifically, we first project the input point cloud into 2D space to generate sparse multi-view images, and then inpaint empty pixels utilizing a pre-trained 2D diffusion model. Next, we design a novel Non-Border-First strategy to unproject the colors of the inpainted dense images back to 3D space, thus obtaining the final textured mesh. In this way, our PointDreamer works in a zero-shot manner, requiring no extra training. Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets show the SoTA performance of PointDreamer, by significantly outperforming baseline methods with 30% improvement in LPIPS score (from 0.118 to 0.068). Code at: https://github.com/YuQiao0303/PointDreamer.
3D Textured Mesh, Colored Point Cloud Reconstruction, Diffusion Models, Image Inpainting
submissionid: ccs: Computing methodologies Reconstructionccs: Computing methodologies 3D imagingccs: Computing methodologies Shape representationsccs: Computing methodologies Appearance and texture representationsccs: Computing methodologies Shape inference
Figure 1. PointDreamer is a zero-shot framework to reconstruct 3D textured meshes from colored point clouds. The key idea is to utilize 2D diffusion priors to inpaint the projected 2D multi-view sparse images to dense ones, which are finally unprojected back to 3D to produce mesh textures. PointDreamer produces textures with enhanced clarity compared to various competitors, such as Screened Poisson Reconstruction (SPR) (Kazhdan and Hoppe, 2013), Neural Kernel Surface Reconstruction (NKSR) (Huang et al., 2023), and Texture Field (Oechsle et al., 2019).
[…]
teaser
1. Introduction
3D textured mesh reconstruction from colored point cloud is a pivotal task in computer vision, graphics and robotics. It enables the transition from sparse and unstructured point clouds to connected and visually-appealing meshes, facilitating applications like digital twin (Tao et al., 2022), metaverse (Wang et al., 2022a), cultural heritage preservation (Wang et al., 2024), etc.
With the great success achieved in geometry reconstruction by learning a 3D implicit field like signed distance field (SDF)  (Park et al., 2019), occupancy field (Mescheder et al., 2019; Chen and Zhang, 2019), etc., researchers propose to learn a 3D texture field (Oechsle et al., 2019) representing color information in 3D space. Nevertheless, such approaches often yield blurry appearances (Tang et al., 2023a). In addition, they necessitate 3D data for training, which is notoriously challenging to acquire, and the limited training data further hinder the models’ generalization capability to unseen objects.
While the 3D textured mesh reconstruction task remains challenging, 2D vision has recently witnessed great prosperity in the field of image super-resolution and image generation, particularly thanks to the emergence of diffusion models (Ho et al., 2020). Compared to 3D vision, 2D vision is more mature, and the abundance of 2D data offers a robust prior, which has been proven crucial for enhancing the performance and generalization ability of deep learning models (Sarker, 2021; Naveed et al., 2023). Inspired by this, numerous studies (Poole et al., 2022; Liu et al., 2023c, d; Tang et al., 2023b; Lin et al., 2023; Qian et al., 2024) seek to adopt 2D diffusion models as a prior to guide 3D generation , mostly by optimizing a score distillation sampling (SDS) (Poole et al., 2022) loss or reconstruction loss between the generated 3D models’ rendered images and 2D diffusion models’ output images. However, they focus on text/image/un-conditioned 3D generation, leaving point-cloud-based reconstruction unexplored. Besides, the colors or textures are still learned or optimized in 3D space (as a texture field or radiance field), yielding blurry appearances (Tang et al., 2023a).
Going beyond existing approaches, our intuition is that, reconstructing 3D textured meshes from colored point clouds is analogous to inpainting 2D sparse images by filling empty pixels. Both tasks focus on somehow dreaming the missing areas of the sparse input, to achieve a more complete and coherent representation. Intuitively, dreaming in 2D space is much easier than in 3D space, especially with powerful tools like diffusion models. This inspires us to convert the dreaming process into 2D space to enhance reconstruction quality especially clarity, and then reconvert the results back to 3D. Specifically, unlike existing works (Gao et al., 2022; Sun et al., 2023; Liu et al., 2024; Poole et al., 2022) that learns per-point colors as implicit functions in 3D space, we predict per-pixel colors in 2D image space by 2D image inpainting, and then directly unproject the high-quality results back to 3D space, so as to generate clear textures for meshes.
Guided by this idea, we design PointDreamer, a novel zero-shot framework for reconstructing high-quality textured meshes from colored point clouds, as shown in Figure 1. With only a set of pre-defined camera parameters from different viewpoints, an off-the-shelf 2D diffusion model, and a geometry extraction module, our method starts from projecting the input point cloud into 2D to generate multi-view sparse images. The 2D diffusion model is then adopted to inpaint the sparse images into dense ones. Finally, we unproject the colors of dense images back to obtain the 3D textured mesh by a carefully-designed Non-Border-First unprojection strategy. It prioritizes non-border areas during unprojection, to circumvent inaccurate color prediction at occlusion borders and thus avoid artifacts. Overall, we list our contributions as follows:
  • We propose PointDreamer, a zero-shot framework for 3D textured mesh reconstruction from colored point cloud. By novelly transforming this 3D reconstruction task into 2D inpainting, PointDreamer achieves high-fidelity appearance and strong generalization ability while requiring no extra training, and pioneers in utilizing 2D diffusion models for 3D point cloud reconstruction.
  • We design Non-Border First (NBF) Unprojection, a novel unprojection strategy for 2D-to-3D re-conversion that effectively avoids artifacts caused by inaccurate color prediction near occlusion borders.
  • Experiments on various benchmarks show the SoTA performance of PointDreamer, by significantly outperforming baseline methods both quantitatively and qualitatively.
2. Related Work
2.1. Textured Mesh Reconstruction from Colored Point Cloud
While many works (Mescheder et al., 2019; Peng et al., 2020; Tang et al., 2021; Boulch and Marlet, 2022) focus on reconstructing surfaces from non-colored point clouds, textured mesh reconstruction from colored point cloud remains a field to be further explored. Traditionally, (Screened) Poisson Surface Reconstruction (SPR) (Kazhdan et al., 2006; Kazhdan and Hoppe, 2013) reconstructs textured meshes by solving the Poisson equation, but it relies on accurate per-point normals as input. In contrast, DHSP3D (Wei et al., 2021) optimizes a MeshCNN in 3D space and XYZ/RGB maps in 2D UV space by self-supervision, but it does not support arbitrary topology. Recently, a more common practice (Gupta et al., 2023; Huang et al., 2023) is to learn two 3D implicit fields or hyper-representations for geometry and texture, respectively. For geometry reconstruction, Signed Distance Field (SDF) (Gupta et al., 2023), Occupancy Field (Chen et al., 2022), Neural Kernel Field (Huang et al., 2023), Deep Marching Tetrahedra (DMTet) (Shen et al., 2021; Gupta et al., 2023), etc. are optional shape representations to be learned or optimized. For texture reconstruction, a texture field (Oechsle et al., 2019) consisting of separated functions for red, green and blue channels is generally fitted, supervised by MAE loss for per-point color (Chen et al., 2022; Huang et al., 2023) or reconstruction loss (Gupta et al., 2023) for images from differentiable rendering. Unfortunately, such methods without powerful prior often suffer from relatively blurry texture (Kerbl et al., 2023). To address this, we are the first, to the best of our knowledge, to perform the 3D reconstruction task by 2D inpainting to take advantage of existing high-performance 2D vision techniques.
2.2. Representations of Colored 3D Objects
Researchers have explored various ways to represent 3D objects or scenes with colors. To begin with, CMR (Kanazawa et al., 2018) and its followers (Li et al., 2020; Goel et al., 2020; Hu et al., 2021; Zhang et al., 2022) regard a shape as the deformation of a sphere with fixed UV mapping, thus only supporting spherical topology. Unlike them, NeRF variants (Mildenhall et al., 2020; Barron et al., 2021; Müller et al., 2022) or DMTet (Shen et al., 2021) with texture field (Oechsle et al., 2019) have shown their advantages in representing power and generation quality  (Gao et al., 2022; Poole et al., 2022; Liu et al., 2023c, d; Tang et al., 2023c; Szymanowicz et al., 2023; Liu et al., 2024; Chen et al., 2023) thanks to the techniques of volumetric rendering and differentiable rendering (Laine et al., 2020). Recently, 3D Gaussian Splatting (Kerbl et al., 2023) has drawn considerable attention where 3D scenes are represented as a set of 3D Gaussians. However, as mentioned before, learning or optimizing colors in 3D space often results in blurry textures (Tang et al., 2023a). In contrast, we directly predict high-fidelity colors in 2D image space by utilizing the power of a 2D pre-trained diffusion model, which can be further unprojected to 3D meshes to obtain clearer textures.
2.3. Diffusion Models for 3D Shape Generation
With the success of 2D diffusion models, researchers have explored extending applications of diffusion models to 3D field. An intuitive idea is to apply the diffusion process on 3D representations, such as a DMTet grid (Liu et al., 2023a), occupancy or SDF field (Zheng et al., 2023), or 3D feature tri-plane (Gupta et al., 2023). To fully capitalize on the abundant availability of 2D data, more researchers seek to directly “lift” results of 2D diffusion models to 3D. Most of these works (Gao et al., 2022; Poole et al., 2022; Lin et al., 2023; Qian et al., 2024; Melas-Kyriazi et al., 2023; Liu et al., 2023c; Shi et al., 2023; Liu et al., 2023d, b; Tang et al., 2023c, b; Szymanowicz et al., 2023; Liu et al., 2024; Chen et al., 2023) optimize a NeRF, a DMTet or 3D Gaussianss by SDS (Poole et al., 2022) or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model (e.g. Stable-Diffusion). View-conditioned diffusion models (Liu et al., 2023c) or multi-view diffusion models (Szymanowicz et al., 2023; Liu et al., 2024) are developed by fine-tuning existing models for generating multi-view-consistent images. Most of these works still learn or optimize colors in 3D space. They tend to be time-consuming during optimization, and suffer from relatively blurry results in unseen views. Instead of optimizing a 3D representation from scratch to follow the images generated by 2D diffusion models, we propose to directly adopt colors of the 2D images, by unprojecting them back to 3D, for higher-quality texture reconstruction.
3. Method
Given a 3D point cloud with per-point XYZ coordinates and RGB colors, our goal is to reconstruct its associated textured mesh. Unlike existing works that learn or optimize colors in 3D space (Oechsle et al., 2019; Poole et al., 2022) or UV space (Wei et al., 2021; Chen et al., 2022), we propose PointDreamer, a novel approach by conducting color prediction in 2D image space to leverage the powerful 2D diffusion priors. Figure 2 presents our designed pipeline with four steps. First, we employ a geometry extraction module to reconstruct an untextured mesh from the input point cloud. Next, with a set of fixed viewpoints, we perform 3D-to-2D conversion by projecting the input point cloud into 2D, producing sparse multi-view images. After that, we conduct color prediction in 2D space by inpainting the empty pixels in the sparse images to form dense ones based on a pre-trained 2D diffusion model (Dhariwal and Nichol, 2021). Finally, we propose a novel Non-Border-First strategy to convert the 2D results back to 3D by unprojecting the colors in the dense images to the untextured mesh to produce the desired textured mesh. Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training on 3D datasets.
Figure 2. Pipeline of our method. Given a colored point cloud, we first reconstruct its untextured mesh. Then, the colored point cloud is projected to 2D to produce sparse multi-view images, which are further inpainted to form the dense ones. Finally, we unproject the results back to 3D to output the associated textured mesh.
Figure 3. Illustration of hidden point removal, without which the results would be messed up.
Figure 4. Top: Illustration of input and output of the unprojection module, which unprojects colors of 2D multi-view images to a 3D mesh. Middle Left: An naive unprojection implementation. For each 3D point $\mathbf{p}$, from all views where $\mathbf{p}$ is visible, the highest-direction-priority view is chosen to paint $\mathbf{p}$. Example point $\mathbf{p}$’s corresponding pixel in the selected view $a$ is with inaccurate color since it’s near an occlusion border, producing an artifact. Bottom Right: our proposed Non-Border-First Unprojection. Observing the correspondence between occlusion border in 2D image space and visibility in UV space (both marked by sky blue dashed lines), we prioritize non-border-areas in UV space to avoid such artifacts.
3.1. Geometry Extraction: Point to Surface
The first step in PointDreamer is to reconstruct an untextured mesh from the input point cloud. Since color information does not need to be considered in this step, many existing point-to-surface methods (Mescheder et al., 2019; Peng et al., 2020; Shen et al., 2021; Boulch and Marlet, 2022) can be employed. In our implementation, we directly adopt the state-of-the-art POCO (Boulch and Marlet, 2022) considering its high performance. Section 4 provides ablation experiments by using different geometry reconstruction methods.
3.2. Projection and Inpainting
As shown in Figure 2, we design a projection module for 3D-to-2D conversion, followed by an inpainting module for 2D color prediction by inpainting sparse multi-view images into dense ones.
3.2.1. Projection: 3D to 2D
Directly projecting the whole point cloud into 2D would incorporate points from occluded parts that should not be visible from the given view. This would mess up the inpainting results and subsequent mesh textures; see the bottom of Figure 3 as an example. Therefore, before generating the sparse images, we precess the input point cloud by the “Hidden Point Removal” operator (Katz et al., 2007), which works by transforming the input and extracting the points that reside on its convex hull. With pre-set camera parameters of $K$ viewpoints ($K=8$ in our implementation) and associated visible point clouds, we conduct camera transformation to these points to get their corresponding pixel coordinates in 2D image space. These pixels are painted according to their associated 3D points’ colors, producing $K$ sparse images. Comparisons of different $K$ values can be found in our supplementary file.
3.2.2. Inpainting: Sparse to Dense
Any 2D inpainting method that fills empty pixels in input images can serve as our inpainting module. We propose to use the state-or-the-art DDNM (Wang et al., 2022b) based on a pre-trained unconditional 2D diffusion model (Dhariwal and Nichol, 2021) to inpaint the multi-view sparse images into dense ones $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$. In this way, the strong prior provided by the 2D diffusion model can facilitate high-quality inpainting. So far, we have completed color prediction purely in 2D space instead of 3D or UV space.
3.3. Unprojection: 2D to 3D
With color prediction conducted in 2D space, we now need to convert the result back to 3D. Specifically, the goal is to use the extracted untextured mesh and inpainted multi-view posed images, to generate the associated textured mesh. Regarding this, we design a novel approach namely “Non-Border-First Unprojection”.
The top of Figure 4 illustrates the basic idea of the unprojection module. As can be seen, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$ via Xatlas (Young, 2022), following (Gao et al., 2022; Long et al., 2023). A texture atlas is an RGB image containing different charts. Each chart is a continuous area in UV space that corresponds to a continuous segment of $\mathcal{M}$ in 3D space. Each pixel $\mathcal{T}[u,v]$ of a chart corresponds to a surface point $\mathbf{p}$ of $\mathcal{M}$. Therefore, our goal is to assign a color for each chart pixel $\mathcal{T}[u,v]$. Given a set of multi-view posed images $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$, $\mathbf{p}$ can be seen in zero, one, or more images. If $\mathbf{p}$ is visible in $I_{k}$, we denote $I_{k}$’s corresponding pixel as $I_{k}[i_{k},j_{k}]$. See the top of Figure 4 as an example, where a 3D point $\mathbf{p}$ is visible in two views $a$ and $b$, and we mark it with a red circle together with its corresponding multi-view image pixels $I_{a}[i_{a},j_{a}]$, $I_{b}[i_{b},j_{b}]$, and texture atlas pixel $\mathcal{T}[u,v]$.
Point $\mathbf{p}$’s visibility in view $k$ can calculated by whether its depth value is no bigger than the corresponding pixel of the $k$’s depth map. If $\mathbf{p}$ is visible only in one view, the corresponding color of $I_{k}[i_{k},j_{k}]$ can be directly adopted as the color of $\mathbf{p}$ and $\mathcal{T}[u,v]$. If $\mathbf{p}$ can be seen in no view at all, we can set some rules to deal with it. If $\mathbf{p}$ can be seen in more than one view, we can try to fuse the corresponding colors or select one best view. Various unprojection strategies differ in how to fuse views or select the best view. Here we introduce three implementations for unprojection.
3.3.1. Naive Unprojection: Only Use Direction Priority
An intuitive idea to pick the best view for $\mathbf{p}$ or $\mathcal{T}[u,v]$ is by direction priority, as shown in the middle left part of Figure 4. Specifically, for each view $k$ where point $\mathbf{p}$ can be seen, we calculate the direction priority score $s_{\mathbf{p}}(k)$, defined as the cosine similarity of $\mathbf{p}$’s normal direction $\mathbf{n}$ and view $k$’s camera direction $\mathbf{d}_{k}$. The view with the highest direction priority is chosen to paint $\mathbf{p}$ and $\mathcal{T}[u,v]$. For example, in the middle left part of Figure 4, $\mathbf{p}$ is painted to gray according to $I_{a}[i_{a},j_{a}]$ since view $a$ has a higher direction priority. Also, if there are a few points invisible from any view, we can still assign a view by direction priority.
While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the middle left part of Figure 4, where a gray artifact is visible in the reconstructed textured mesh that should have been brown. Through analysis, such artifacts appear in border areas, that is, the area near the border of occluded (e.g. bottom side of the brown seat in Figure 4) and occluding part (e.g. the gray bottom of the chair) of the shape in a posed image $I_{k}$ when the former is occluded by the latter from view $k$ (e.g. view $a$). In such areas, the inpainting module may inaccurately delineate the boundary separating the two parts, e.g. too large gray bottom in view $a$, resulting in artifacts.
3.3.2. Optimization-based Unprojection
Instead of selecting one single view by direction priority, 3D generation method DreamGaussian (Tang et al., 2023a) adopts an optimization-based unprojection strategy to get the textured mesh from multi-view images. The idea is to optimize the texture atlas by the multi-view per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh. In this way, the optimization process kinds of fuses corresponding colors in different views instead of selecting the best one. However, it cannot ignore the inaccurate colors in the border areas, so the artifacts can only be reduced instead of eliminated. This inspires us to design a non-border-first strategy to address this issue.
3.3.3. Non-Border-First Unprojection (NBF)
The bottom of Figure 4 illustrates the idea of our proposed novel unprojection strategy namely Non-Border-First Unprojection, which begins with detecting border areas in multi-view images, so as to paint the texture atlas while ignoring them. After that, if some atlas pixels remain unpainted since their associated 3D points can not be seen from any non-border area, we then select from all areas, whether they are border areas or not, to paint the rest pixels. Now the question is, how to accurately detect border areas.
Interestingly, we find that, the borders between occluded and occluding parts of the shape in image $I_{k}$ (2D space) often correspond to the borders between visible and invisible areas in “view $k$ visible texture atlas $\mathcal{T}_{k}$” (UV space), where $\mathcal{T}_{k}$ is defined as a texture atlas in which only visible areas in view $k$ are filled and rest invisible areas are empty. See the bottom of Figure 4 as an example, where such border areas are marked by sky blue dashed lines. In posed images $I_{a}$ and $I_{b}$, these areas divide the occluded brown seat and the occluding gray bottom, while in view $a$ and $b$ visible texture atlas $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$, these areas divide the visible (painted) and invisible (transparent) parts in view $a$ and $b$, respectively. The reason for this correspondence of UV visibility border and 2D occlusion border is that, a continuous area within a chart of the texture atlas usually corresponds to a continuous segment of the 3D surface. When a chart contains both visible and invisible regions in a given view $k$, it often indicates that the corresponding 3D areas of invisible regions are occluded by some other part of the shape. Therefore, we design our Non-Border-First unprojection strategy as follows.
  1. (1)
    We paint “view $k$ visible texture atlas map” $\mathcal{T}_{k}$ for each view $k$ defined as above.
  2. (2)
    We calculate the border areas by dilating the edges between visible and invisible areas in $\mathcal{T}_{k}$ given a certain dilation kernel size as a hyperparameter.
  3. (3)
    We paint the final texture atlas $\mathcal{T}$ considering only non-border areas in ${\mathcal{T}_{k}}$. During this, if a point is visible in more than one view’s non-border-areas, we select the best view with the highest direction priority.
  4. (4)
    If there remain some unpainted pixels in $\mathcal{T}$, we now check the previously ignored border areas to paint them, during which we still select among views by direction priority.
  5. (5)
    If some points cannot be seen from either view, we assign it with a view by direction priority considering all areas in all views whether they’re border areas or not.
  6. (6)
    We experimentally find that additionally optimizing from the generated texture atlas by only calculating the loss of non-border pixels would further enhance the performance.
In this way, our non-border-first strategy can reasonably eliminate artifacts caused by inaccurate color prediction in border areas. We conduct experiments to verify the effectiveness of our proposed strategy in Section 4.
4. Experiments and Results
4.1. Experimental Settings
4.1.1. Datasets
To evaluate the performance of our method against other competitors, we use three benchmark datasets in our experiments.
(i) ShapeNetCoreV2 (Chang et al., 2015): A large-scale synthesis 3D object dataset, where we follow (Gao et al., 2022) to use the official test splits of chair, car, and motorbike categories for evaluation since they contain relatively complex textures. The test sets of the three categories contain around 1300, 690 and 70 samples, respectively.
(ii) Google Scanned Objects (GSO) (Downs et al., 2022): A real-scanned 3D object dataset and we use all its 1030 samples for evaluation unlike existing works (Liu et al., 2024; Long et al., 2023) that only use 30 of them.
(iii) OmniObject3D (Wu et al., 2023): A real-scanned 3D object dataset with 6000 samples. For efficiency, we randomly select 100 objects for evaluation.
For each textured mesh from the above datasets, we sample 30000 colored points as input. Note that, our approach works in a zero-shot manner, which can be directly applied to objects from various datasets or categories, requiring no extra training but only an off-the-shelf 2D diffusion model (Dhariwal and Nichol, 2021).
4.1.2. Metrics
To evaluate the quality of the reconstructed 3D textured mesh, we follow existing works (Liu et al., 2023c, 2024; Long et al., 2023; Zhang et al., 2022) to compare the similarity between multi-view images rendered by the reconstructed mesh and ground-truth mesh. Specifically, we render 20 multi-view images on each mesh for a thorough assessment and four similarity metrics are calculated: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) (Wang et al., 2004), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), and Fréchet Inception Distance (FID)  (Heusel et al., 2017). PSNR measures pixel-level similarity. SSIM considers both pixel values and structural information. LPIPS uses a pre-trained deep neural network to assess image similarity, imitating human perception better than PSNR and SSIM. FID measures the statistical distance between the distribution of two datasets. PSNR and SSIM scores indicate higher quality as their values increase, whereas lower values of LPIPS and FID reflect better performance.
4.2. Baselines
To fully evaluate the performance of our method, we compare it with three kinds of baselines:
(1) Non-deep-learning method: We choose the classical Screened Poisson Reconstruction (SPR) (Kazhdan and Hoppe, 2013) for comparison. Note that SPR requires per-point normal as input, and we estimate it via Principal Component Analysis (PCA).
(2) DHSP3D (Wei et al., 2021) and NKSR (Huang et al., 2023): These are two recent networks for textured mesh reconstruction and we employ the official code provided by the authors. Similar to SPR, NKSR also requires point normals as additional input.
(3) Texture Field (Oechsle et al., 2019): Since we cannot find any existing open-source texture-field-based method requiring only colored point clouds as input, we thus design a baseline network inspired by (Gupta et al., 2023). The key idea is to learn a 3D feature tri-plane from the input point cloud and then decode RGB values for the query points. We train it on the official training split of the mentioned three categories of ShapeNetCoreV2, supervised by per-point color MSE loss following (Gupta et al., 2023). During inference, we employ the above-mentioned POCO (Boulch and Marlet, 2022) to reconstruct untextured meshes and utilize our trained texture field network to infer textures by querying colors for the associated 3D point of each texture atlas pixel. In this way, both our proposed PointDreamer and the texture field baseline adopt the same geometry module for a fair comparison. Please refer to our supplementary file for more details about the training and inference of our texture field baseline.
Figure 5. Qualitative results of textured mesh reconstruction on ShapeNetCoreV2 dataset.
Figure 6. Qualitative results of textured mesh reconstruction on Google Scanned Objects dataset.
Figure 7. Qualitative results of textured mesh reconstruction on OmniObject3D dataset.
[…]
Figure 8. Qualitative results on textured mesh reconstruction comparing our method and DHSP3D.
Table 1. Comparisons on textured mesh reconstruction performance on ShapeNetCoreV2 dataset.
{NiceTabular}
c—c—cccc ShapeNet Cat. Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
Chairs SPR 24.10 0.931 0.0923 12.03
NKSR 23.05 0.931 0.0964 29.11
T.F. 26.12 0.947 0.0664 6.83
Ours 26.29 0.952 0.0570 4.90
Cars SPR
23.46 0.929 0.0882 43.89
NKSR 21.01 0.919 0.0961 131.05
T.F. 22.42 0.919 0.0868 38.08
Ours 22.78
0.930 0.0727 12.41
Motorbikes SPR 15.09 0.809 0.2139 106.09
NKSR 18.13 0.905 0.1008 175.42
T.F. 21.17 0.926 0.0635 40.06
Ours
21.20 0.930 0.0562 28.42
Table 2. Comparisons on textured mesh reconstruction performance on GSO and Omniobject3D datasets. “T.F” denotes Texture Field.
{NiceTabular}
c—c—cccc Dataset Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
GSO SPR 27.11 0.907 0.1200 27.43
NKSR 24.65 0.900 0.1184 36.51
T.F. 25.53 0.894 0.1276 46.44
Ours 27.24 0.923 0.0827 9.32
OmniObject3D SPR
30.20 0.927 0.0998 40.13
NKSR 26.77 0.919 0.1048 50.14
T.F. 28.08 0.914 0.1121 65.61
Ours 29.78
0.941 0.0683 18.23
4.3. Main Results
4.3.1. Qualitative comparisons
Since DHSP3D is much slower than other methods (about 4-6 hours per shape compared to ¡ 2 min), we randomly select two objects from GSO dataset for efficiency when comparing our method against DHSP3D; see Figure 8. For other methods (i.e., SPR, NKSR & texture field), we present the visual comparisons on each dataset separately; see Figures 56 and 7.
Figures 57 show that our method, though zero-shot, achieves much clearer and more realistic textures than baselines, thanks to the strong 2D diffusion prior. Besides, SPR and NKSR, which both rely on per-point normals as input, sometimes yield redundant geometry mainly due to the wrongly-estimated normal direction. The texture field baseline, which is trained on ShapeNetCoreV2, performs much worse on the newly-seen dataset of GSO and Omniobject3d, indicating a lack of generalization ability.
In Figure 8, the red cup in the first row shows that DHSP3D does not support arbitrary topology. This is because it optimizes meshCNN and 2D XYZ map from the convex hull of the input point cloud, enforcing the output mesh to adhere to the hull’s topology. Besides, DHSP3D suffers from less realistic textures with jagged or unclear edges, since its purely self-prior property only enables it to predict point colors considering corresponding neighbors, without the ability to dream the unseen like our PointDreamer.
4.3.2. Quantitative comparisons
We summarize the quantitative results in Tables 4.2 and 4.2. We can see that our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin. However, regarding PSNR, which prioritizes pixel-level accuracy and thus may differ from human perception of quality, our method is sometimes outperformed by SPR. We assume that SPR predicts a point’s color by somehow fusing nearby points’ colors, leading to blurry textures but safely avoiding extreme pixel-level errors. On the contrary, our method dreams the colors by its diffusion prior. While achieving visually better results, a few pixels with extremely high errors may lower the overall PSNR score.
4.4. Impact of Degraded Input Quality
4.4.1. Anti-Noise Ability Analysis
We compare the anti-noise abilities of our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset. Table 4.5 shows the results, where our method, even with noisy input, outperforms baselines with clean input, indicating a strong anti-noise ability. Visual comparisons can be found in our supplementary file.
4.4.2. Sparsity Test
To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 20k gradually, and present the results in Table 4.5. We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input; see results in Table 4.2 “Chairs” category. Again, visual comparisons can be found in our supplementary file.
4.5. Method Component Analysis
We here compare different implementations for sub-modules (geometry extraction, inpainting and unprojection module) of our method by replacing one module at a time while keeping others unchanged. All experiments in this subsection are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
Table 3. Comparisons on anti-noise ability on chair category of ShapeNetCoreV2 dataset. “Noisy” denotes adding Gaussian noise with a standard deviation of 0.005 to the input point cloud.
{NiceTabular}
c—cc—cc—cc—cc
Method PSNR ↑   SSIM ↑   LPIPS ↓   FID ↓
Clean Noisy Clean Noisy Clean Noisy Clean Noisy
SPR 24.10 19.64 0.9311 0.8840 0.0923 0.1702 12.03 65.25
NKSR 23.05 22.79 0.9307 0.9288 0.0964 0.1017 29.11 44.02
T.F. 26.12 26.01 0.9467 0.9436 0.0664 0.0711 6.83 8.21
Ours 26.29 26.26 0.9524 0.9516 0.0570 0.0574 4.90 4.93
Table 4. Comparisons on our method’s performance with increased input point cloud sparsity.
{NiceTabular}
c—cccc Point Number PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
30k 26.2910 0.9524 0.0570 4.9048
25k 26.2761 0.9517 0.0580 5.2412
20k 26.2553 0.9509 0.0594 5.6960
4.5.1. Geometry Extraction
In addition to POCO, we further explore extracting geometry by SPR and depth inpainting to investigate how the geometry extraction module influence the final reconstructed textured mesh. Depth inpainting refers to inpainting projected sparse 2D depth maps instead of RGB images, and then reconstructing geometry from the inpainted dense depth maps by depth fusion (Griegler, 2017). Please refer to our supplementary file for more details and the quantitative results. We present the visual comparisons in Figure 9. We can see that, compared to POCO, both SPR and depth inpainting suffer from noisy geometry, resulting in noisy textures. Clearly, a higher-performance geometry extraction module contributes to a more refined reconstructed textured mesh.
Figure 9. Geometry extraction module replacement results comparing our POCO against SPR and Depth Inpainting.
4.5.2. 2D Inpainting
We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR (Zhu et al., 2023) and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability. We provide the visual comparisons in Figure 10 and quantitative results in our supplementary file. As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, indicating the importance of a strong inpainting module to the reconstruction performance.
Figure 10. Inpainting module replacement results comparing our adopted DDNM against nearest interpolation, linear interpolation, and DiffPIR.
4.5.3. Unprojection
We present the comparison results of different unprojection strategies in Figure  11 and Table 4.5.3.
  • “Opt. Scratch” denotes optimizing the texture atlas from scratch (Tang et al., 2023a);
  • “Naive” selects the best views by direction priority;
  • “NBF” selects the best views by our proposed Non-Border-First strategy;
  • “Opt. Naive” further optimizes the obtained atlas of “Naive”;
  • “Opt. NBF” further optimizes the obtained atlas of “NBF” while considering only loss in non-border areas.
The first row of Figure 11 shows an example, where optimizing from scratch fails on areas unseen from any view due to occlusion. Both rows show that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization. Only our NBF and Opt. NBF nearly eliminate such artifacts. Table 4.5.3 also shows that our proposed NBF and Opt. NBF outperform other unprojection strategies.
Figure 11. Qualitative results of replacing our unprojection strategy with others. “Opt.” denotes “Optimize”.
Table 5. Quantitative results of replacing our NBF unprojection strategy with others. “Opt.” denotes “Optimize”.
{NiceTabular}
c—cccc Unprojection Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
Opt. Scratch (Tang et al., 2023a) 26.1939 0.9473 0.0592 5.2619
Naive 26.2225 0.9500 0.0587 5.1068
Opt. Naive 26.2477 0.9511 0.0580 4.9389
NBF 26.2729 0.9512 0.0579 5.0309
Opt. NBF 26.2565
0.9516 0.0574 4.9326
5. Conclusion, Limitations and Future Work
Conclusion. We propose PointDreamer, a novel zero-shot framework for textured mesh reconstruction from colored point cloud. PointDreamer predicts colors by 2D inpainting to utilize the strong diffusion prior and achieves SoTA performance. We also propose a novel “Non-Border-First” strategy to unproject the colors of predicted 2D images back to 3D space.
Limitation and Future Work. The main shortcoming of our proposed method is its relatively low speed. We employed one NVIDIA GeForce GTX 3090 GPU and it takes about 100s for our method to reconstruct a textured mesh, compared to NKSR’s 0.64 s. It takes about 72 s for DDNM (Wang et al., 2022b) to inpaint our 8 multi-view images, which takes most of the time. Replacing DDNM with a faster inpainting module may increase our reconstruction speed. In the future, we may seek to address this issue, and further explore 3D-aware inpainting and relighting-supported texture reconstruction.
Appendix A Appendix: Implementation Details
A.1. Texture Field Baseline
A.1.1. Network architecture
Inspired by works (Gupta et al., 2023; Shen et al., 2021) that represent 3D information by a feature tri-plane, we adopt the network architecture of the open-source Convolutional Occupancy Network (Peng et al., 2020), which also follows a tri-plane representation. Specifically, we modify its one-dimensional output head for occupancy prediction to three-dimensional for RGB color prediction.
A.1.2. Training
We follow 3DGen (Gupta et al., 2023) to train the network by the per-3D-point MSE loss on predicted and GT colors of sampled 3D points. Additionally, we also tried to employ the per-2D-pixel MSE loss of 2D images rendered from GT meshes and predicted meshes by differentiable rendering, but the experimental results are worse, as shown in Figure 12. Therefore, we adopt the per-3D-point MSE loss in our manuscript. We train the texture field network on a single NVIDIA GeForce GTX 3090 GPU with a batch size of 24 and a learning rate of 0.0001 for 2592, 024 iterations (about 356 epochs).
A.1.3. Inference.
During inference, we first use POCO (Boulch and Marlet, 2022) to predict an untextured mesh from the input point cloud, and apply UV mapping to it by Xatlas (Young, 2022), which produces the 3D positions of each valid pixel in the texture atlas. We query the color of each of these 3D positions by our trained texture field network, to inpaint the texture atlas, so as to obtain the final textured mesh.
Figure 12. Visual comparisons of the Texture Field baseline trained with different losses. “2D per-pixel” denotes rendering the generated textured mesh to multi-view 2D images, and then calculating the MSE loss between the rendered and GT images. “3D per-point” denotes calculating MSE loss between the predicted and GT colors of sampled 3D points.
A.2. Depth Inpainting for Geometry Extraction
With 2D inpainting adopted as the key for our texture reconstruction, we can also introduce it to geometry reconstruction, by inpainting 2D depth maps instead of RGB images. Specifically, we follow the following steps to reconstruct untextured meshes from the input point cloud, as shown in Figure 13:
Figure 13. Pipeline of extracting geometry by depth inpainting.
  1. (1)
    We generate multi-view sparse depth maps by projecting 3D points to 2D, and assigning the value of each pixel as the depth value of the corresponding 3D point. Note that, similar to generating our sparse RGB images, hidden point removal is conducted for each viewpoint before projecting.
  2. (2)
    Since depth fusion requires depth maps’ background pixels to have infinite values to produce a reasonable mesh, we generate foreground masks by projecting 3D points to 2D space with a relatively big point size, i.e. the number of 2D pixels occupied by each 3D point. In this way, most foreground pixels (pixels that should correspond to a point on the 3D mesh) can be occupied, and we use the closing operation of morphology to fill the rest small holes. In addition, since we use a big point size to generate the foreground mask, the mask would be bigger than the ground truth, so we shrink the generated mask by erosion.
  3. (3)
    We inpaint the foreground pixels of the sparse depth maps into dense ones by nearest interpolation considering efficiency.
  4. (4)
    We produce an untextured mesh by depth fusion based on the inpainted dense depth maps, and conduct mesh simplification (Garland and Heckbert, 1997) and Taubin Smooth (Taubin, 1995) to it as post-processing, to get the final untextured mesh.
Appendix B Appendix: More Experimental Results
B.1. Effect of Different $K$ Values (Number of Viewpoints for Projection and Inpainting)
To investigate the effect of different numbers of viewpoints (denoted as $K$) for projection and inpainting, we conduct experiments on the motorbike category or ShapeNetCoreV2 dataset by using different $K$ values including 6, 8, and 20. Table B.1 shows the distribution of cameras for each setting, together with the quantitative results. We can see that more views contribute to a slightly higher reconstruction quality. Visual comparisons in Figure 14 also show that, an insufficient number of views would produce artifacts in invisible or occluded areas, thus impacting the performance.
Considering that using $K=20$ views is only slightly better than setting $K=8$, but inpainting more views’ images can be much more time-consuming, we set the number of views to be 8 for most datasets in our manuscript to balance both effectiveness and efficiency. The only exception is the motorbike category from the ShapeNetCoreV2 dataset, for which we use the 20 views instead, considering motorbikes’ more complex geometry and topology.
Table 6. Quantitative results of using different $K$ values (numbers of viewpoints for projection and inpainting) on the motorbike category of ShapeNetCoreV2 dataset. More views contribute to a slightly higher reconstruction quality.
{NiceTabular}
cc—cccc Viewpoint Number Camera Distribution   PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
6 At the centers of each face of a cube 20.8755 0.9273 0.0585 33.2870
8 Evenly distributed on a Fibonacci Sphere 21.0664 0.9287 0.0572 30.2113
20 On the 20 vertices of a regular icosahedron 21.2013 0.9299 0.0562 28.4213
Figure 14. Visual comparisons of our reconstructed meshes with different $K$ values (numbers of viewpoints for projection and inpainting). An insufficient number of views would lead to artifacts in invisible or occluded areas.
B.2. Visual Comparisons: Impact of Degraded Input Quality
B.2.1. Anti-Noise Ability Analysis
Figure 15 presents the visual comparisons of our method and baseline methods’ reconstructed meshes from noisy or clean input point clouds. Overall, we have the following observations, which are consistent with the quantitative results in our manuscript:
  1. (1)
    As expected, all methods produce higher-quality textured meshes with clean input point clouds compared to noisy ones.
  2. (2)
    Our PointDreamer shows a relatively high anti-noise ability, where only a small performance drop is observed when giving noisy input.
  3. (3)
    Our PointDreamer with noisy input point clouds shows an even better visual effect compared to baseline methods with clean inputs.
Figure 15. Visual comparisons of our PointDreamer’s and baseline methods’ reconstructed meshes, with noisy or clean point clouds as input. Our PointDreamer shows a strong anti-noise ability by producing high-quality textures even with noisy input.
B.2.2. Sparsity Test
Figure 16 shows the visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input. There is a very small performance drop introduced by decreasing the input point number, which can sometimes be hard to notice by human eyes. This indicates a relatively high robustness of our method towards varying degrees of input sparsity.
Figure 16. Visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input. There is a small performance drop when adopting sparser input, which sometimes can be hard to notice by human eyes.
B.3. Quantitative Results of Method Component Analysis
B.3.1. Geometry Extraction
We provide the quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting in Table B.3.1. As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID. This is consistent with the visual comparisons (see Figure 9) in our manuscript where both SPR and Depth Inpainting suffer from noisy geometry and thus low-quality textures.
Table 7. Quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting.
{NiceTabular}
c—cccc Geometry Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
SPR 19.4071 0.8752 0.1777 74.9854
Depth Inpainting 26.2502 0.9466 0.0717 20.7198
Our Adopted POCO 26.2565 0.9516 0.0574 4.9326
B.3.2. 2D Inpainting
We provide the quantitative results of replacing our 2D inpainting module from DDNM (Wang et al., 2022b) to nearest interpolation, linear interpolation, and DiffPIR (Zhu et al., 2023) in Table B.3.2. DiffPIR is another 2D-diffusion-based image restoration approach with inpainting ability. As can be seen, DDNM outperforms the other three methods thanks to its strong inpainting ability, which is consistent with the visual comparisons (see Figure 10) in our manuscript.
Table 8. Quantitative results of replacing our inpainting module from DDNM to nearest interpolation, linear interpolation, and DiffPIR.
{NiceTabular}
c—cccc Inpainting Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
Nearest Interpolation 26.1175 0.9457 0.0618 11.0205
Linear Interpolation 26.1739 0.9474 0.0612 9.4725
DiffPIR 26.1582 0.9456 0.0652 9.4823
Our Adopted DDNM 26.2565 0.9516 0.0574 4.9326

### v2

@@
-PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud by 2D Inpainting
-Abstract.
-Reconstructing textured meshes from colored point clouds is an important but challenging task in 3D graphics and vision.
-Most existing methods predict colors as implicit functions in 3D or UV space, suffering from blurry textures or the lack of generalization capability.
-Addressing this, we propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud.
-It produces meshes with enhanced fidelity and clarity by 2D image inpainting, taking advantage of the mature techniques and massive data of 2D vision.
-Specifically, we first project the input point cloud into 2D space to generate sparse multi-view images, and then inpaint empty pixels utilizing a pre-trained 2D diffusion model.
-Next, we design a novel Non-Border-First strategy to unproject the colors of the inpainted dense images back to 3D space, thus obtaining the final textured mesh.
-In this way, our PointDreamer works in a zero-shot manner, requiring no extra training.
+PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud
+Abstract
+Reconstructing textured meshes from colored point clouds is an important but challenging task.
+Most existing methods yield blurry-looking textures or rely on 3D training data that are hard to acquire.
+Regarding this, we propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud via diffusion-based 2D inpainting.
+Specifically, we first reconstruct an untextured mesh.
+Next, we project the input point cloud into 2D space to generate sparse multi-view images, and then inpaint empty pixels utilizing a pre-trained 2D diffusion model.
+After that, we unproject the colors of the inpainted dense images onto the untextured mesh, thus obtaining the final textured mesh.
+This project-inpaint-unproject pipeline bridges the gap between 3D point clouds and 2D diffusion models for the first time.
+Thanks to the powerful 2D diffusion model pre-trained on extensive 2D data, PointDreamer reconstructs clear, high-quality textures with high robustness to sparse or noisy input.
+Also, it’s zero-shot requiring no extra training.
+In addition, we design Non-Border-First unprojection strategy to address the border-area inconsistency issue, which is less explored but commonly-occurred in methods that generate 3D textures from multiview images.
 Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets show the SoTA performance of PointDreamer, by significantly outperforming baseline methods with 30% improvement in LPIPS score (from 0.118 to 0.068).
 Code at: https://github.com/YuQiao0303/PointDreamer.
-3D Textured Mesh, Colored Point Cloud Reconstruction, Diffusion Models, Image Inpainting
-submissionid: ccs: Computing methodologies Reconstructionccs: Computing methodologies 3D imagingccs: Computing methodologies Shape representationsccs: Computing methodologies Appearance and texture representationsccs: Computing methodologies Shape inference
-Figure 1.
-PointDreamer is a zero-shot framework to reconstruct 3D textured meshes from colored point clouds.
-The key idea is to utilize 2D diffusion priors to inpaint the projected 2D multi-view sparse images to dense ones, which are finally unprojected back to 3D to produce mesh textures.
-PointDreamer produces textures with enhanced clarity compared to various competitors, such as Screened Poisson Reconstruction (SPR) (Kazhdan and Hoppe, 2013), Neural Kernel Surface Reconstruction (NKSR) (Huang et al., 2023), and Texture Field (Oechsle et al., 2019).
-[…]
-teaser
-1. Introduction
-3D textured mesh reconstruction from colored point cloud is a pivotal task in computer vision, graphics and robotics.
-It enables the transition from sparse and unstructured point clouds to connected and visually-appealing meshes, facilitating applications like digital twin (Tao et al., 2022), metaverse (Wang et al., 2022a), cultural heritage preservation (Wang et al., 2024), etc.
-With the great success achieved in geometry reconstruction by learning a 3D implicit field like signed distance field (SDF)  (Park et al., 2019), occupancy field (Mescheder et al., 2019; Chen and Zhang, 2019), etc., researchers propose to learn a 3D texture field (Oechsle et al., 2019) representing color information in 3D space.
-Nevertheless, such approaches often yield blurry appearances (Tang et al., 2023a).
-In addition, they necessitate 3D data for training, which is notoriously challenging to acquire, and the limited training data further hinder the models’ generalization capability to unseen objects.
-While the 3D textured mesh reconstruction task remains challenging, 2D vision has recently witnessed great prosperity in the field of image super-resolution and image generation, particularly thanks to the emergence of diffusion models (Ho et al., 2020).
-Compared to 3D vision, 2D vision is more mature, and the abundance of 2D data offers a robust prior, which has been proven crucial for enhancing the performance and generalization ability of deep learning models (Sarker, 2021; Naveed et al., 2023).
-Inspired by this, numerous studies (Poole et al., 2022; Liu et al., 2023c, d; Tang et al., 2023b; Lin et al., 2023; Qian et al., 2024) seek to adopt 2D diffusion models as a prior to guide 3D generation , mostly by optimizing a score distillation sampling (SDS) (Poole et al., 2022) loss or reconstruction loss between the generated 3D models’ rendered images and 2D diffusion models’ output images.
-However, they focus on text/image/un-conditioned 3D generation, leaving point-cloud-based reconstruction unexplored.
-Besides, the colors or textures are still learned or optimized in 3D space (as a texture field or radiance field), yielding blurry appearances (Tang et al., 2023a).
-Going beyond existing approaches, our intuition is that, reconstructing 3D textured meshes from colored point clouds is analogous to inpainting 2D sparse images by filling empty pixels.
-Both tasks focus on somehow dreaming the missing areas of the sparse input, to achieve a more complete and coherent representation.
-Intuitively, dreaming in 2D space is much easier than in 3D space, especially with powerful tools like diffusion models.
-This inspires us to convert the dreaming process into 2D space to enhance reconstruction quality especially clarity, and then reconvert the results back to 3D.
-Specifically, unlike existing works (Gao et al., 2022; Sun et al., 2023; Liu et al., 2024; Poole et al., 2022) that learns per-point colors as implicit functions in 3D space, we predict per-pixel colors in 2D image space by 2D image inpainting, and then directly unproject the high-quality results back to 3D space, so as to generate clear textures for meshes.
-Guided by this idea, we design PointDreamer, a novel zero-shot framework for reconstructing high-quality textured meshes from colored point clouds, as shown in Figure 1.
-With only a set of pre-defined camera parameters from different viewpoints, an off-the-shelf 2D diffusion model, and a geometry extraction module, our method starts from projecting the input point cloud into 2D to generate multi-view sparse images.
-The 2D diffusion model is then adopted to inpaint the sparse images into dense ones.
-Finally, we unproject the colors of dense images back to obtain the 3D textured mesh by a carefully-designed Non-Border-First unprojection strategy.
-It prioritizes non-border areas during unprojection, to circumvent inaccurate color prediction at occlusion borders and thus avoid artifacts.
-Overall, we list our contributions as follows:
-  • We propose PointDreamer, a zero-shot framework for 3D textured mesh reconstruction from colored point cloud.
-By novelly transforming this 3D reconstruction task into 2D inpainting, PointDreamer achieves high-fidelity appearance and strong generalization ability while requiring no extra training, and pioneers in utilizing 2D diffusion models for 3D point cloud reconstruction.
-  • We design Non-Border First (NBF) Unprojection, a novel unprojection strategy for 2D-to-3D re-conversion that effectively avoids artifacts caused by inaccurate color prediction near occlusion borders.
-  • Experiments on various benchmarks show the SoTA performance of PointDreamer, by significantly outperforming baseline methods both quantitatively and qualitatively.
-2. Related Work
-2.1. Textured Mesh Reconstruction from Colored Point Cloud
-While many works (Mescheder et al., 2019; Peng et al., 2020; Tang et al., 2021; Boulch and Marlet, 2022) focus on reconstructing surfaces from non-colored point clouds, textured mesh reconstruction from colored point cloud remains a field to be further explored.
-Traditionally, (Screened) Poisson Surface Reconstruction (SPR) (Kazhdan et al., 2006; Kazhdan and Hoppe, 2013) reconstructs textured meshes by solving the Poisson equation, but it relies on accurate per-point normals as input.
-In contrast, DHSP3D (Wei et al., 2021) optimizes a MeshCNN in 3D space and XYZ/RGB maps in 2D UV space by self-supervision, but it does not support arbitrary topology.
-Recently, a more common practice (Gupta et al., 2023; Huang et al., 2023) is to learn two 3D implicit fields or hyper-representations for geometry and texture, respectively.
-For geometry reconstruction, Signed Distance Field (SDF) (Gupta et al., 2023), Occupancy Field (Chen et al., 2022), Neural Kernel Field (Huang et al., 2023), Deep Marching Tetrahedra (DMTet) (Shen et al., 2021; Gupta et al., 2023), etc.
-are optional shape representations to be learned or optimized.
-For texture reconstruction, a texture field (Oechsle et al., 2019) consisting of separated functions for red, green and blue channels is generally fitted, supervised by MAE loss for per-point color (Chen et al., 2022; Huang et al., 2023) or reconstruction loss (Gupta et al., 2023) for images from differentiable rendering.
-Unfortunately, such methods without powerful prior often suffer from relatively blurry texture (Kerbl et al., 2023).
-To address this, we are the first, to the best of our knowledge, to perform the 3D reconstruction task by 2D inpainting to take advantage of existing high-performance 2D vision techniques.
-2.2. Representations of Colored 3D Objects
-Researchers have explored various ways to represent 3D objects or scenes with colors.
-To begin with, CMR (Kanazawa et al., 2018) and its followers (Li et al., 2020; Goel et al., 2020; Hu et al., 2021; Zhang et al., 2022) regard a shape as the deformation of a sphere with fixed UV mapping, thus only supporting spherical topology.
-Unlike them, NeRF variants (Mildenhall et al., 2020; Barron et al., 2021; Müller et al., 2022) or DMTet (Shen et al., 2021) with texture field (Oechsle et al., 2019) have shown their advantages in representing power and generation quality  (Gao et al., 2022; Poole et al., 2022; Liu et al., 2023c, d; Tang et al., 2023c; Szymanowicz et al., 2023; Liu et al., 2024; Chen et al., 2023) thanks to the techniques of volumetric rendering and differentiable rendering (Laine et al., 2020).
-Recently, 3D Gaussian Splatting (Kerbl et al., 2023) has drawn considerable attention where 3D scenes are represented as a set of 3D Gaussians.
-However, as mentioned before, learning or optimizing colors in 3D space often results in blurry textures (Tang et al., 2023a).
-In contrast, we directly predict high-fidelity colors in 2D image space by utilizing the power of a 2D pre-trained diffusion model, which can be further unprojected to 3D meshes to obtain clearer textures.
-2.3. Diffusion Models for 3D Shape Generation
-With the success of 2D diffusion models, researchers have explored extending applications of diffusion models to 3D field.
-An intuitive idea is to apply the diffusion process on 3D representations, such as a DMTet grid (Liu et al., 2023a), occupancy or SDF field (Zheng et al., 2023), or 3D feature tri-plane (Gupta et al., 2023).
-To fully capitalize on the abundant availability of 2D data, more researchers seek to directly “lift” results of 2D diffusion models to 3D.
-Most of these works (Gao et al., 2022; Poole et al., 2022; Lin et al., 2023; Qian et al., 2024; Melas-Kyriazi et al., 2023; Liu et al., 2023c; Shi et al., 2023; Liu et al., 2023d, b; Tang et al., 2023c, b; Szymanowicz et al., 2023; Liu et al., 2024; Chen et al., 2023) optimize a NeRF, a DMTet or 3D Gaussianss by SDS (Poole et al., 2022) or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model (e.g.
-Stable-Diffusion).
-View-conditioned diffusion models (Liu et al., 2023c) or multi-view diffusion models (Szymanowicz et al., 2023; Liu et al., 2024) are developed by fine-tuning existing models for generating multi-view-consistent images.
-Most of these works still learn or optimize colors in 3D space.
-They tend to be time-consuming during optimization, and suffer from relatively blurry results in unseen views.
-Instead of optimizing a 3D representation from scratch to follow the images generated by 2D diffusion models, we propose to directly adopt colors of the 2D images, by unprojecting them back to 3D, for higher-quality texture reconstruction.
-3. Method
+Index Terms:
+Point Cloud Reconstruction, Texture Mapping, 2D Diffusion, Image Inpainting.
+1 Introduction
+3D point clouds and meshes are two of the most fundamental 3D modalities in computer vision, graphics, and robotics.
+Reconstructing 3D textured meshes from colored point clouds (abbr.
+colored-PC-to-mesh) is a pivotal task, with extensive applications like digital twin [1], metaverse [2], cultural heritage preservation [3], etc.
+It requires the accurate reconstruction of both geometry and colors, to transform sparse and unstructured data into coherent and visually appealing models.
+While recent methods [4, 5, 6] have achieved relatively satisfactory geometry reconstruction quality, color reconstruction remains challenging, especially for low-density point clouds.
+Unlike images that provide dense pixel-level representations, the inherent sparse nature of point clouds makes it difficult to reconstruct dense colors.
+Regarding this, existing methods either blend input points’ colors by interpolation [7] or overfitting [8, 9], or train color prediction networks [10, 11] with 3D datasets.
+However, they often yield blurring-looking textures; see SPR [7], NKSR [9], and Texture Field [12, 10] in Figure 1.
+Also, 3D training data are notoriously challenging to acquire.
+Figure 1: PointDreamer is a zero-shot framework to reconstruct 3D textured meshes from colored point clouds.
+It produces textures with higher quality compared to baseline methods.
+Its specially designed NBF unprojection strategy effectively addresses the border-area artifact issue.
+While the colored-PC-to-mesh task remains challenging, 2D image generation has recently thrived with impressive plausibility, level of detail, and generalization capability.
+Therefore, many works [13, 14] leverage 2D diffusion models for 3D generation.
+Since the scale of existing 2D datasets far exceeds that of 3D datasets, leveraging diffusion models pre-trained on extensive 2D datasets can also significantly alleviate the reliance on 3D datasets.
+This inspires us to adopt 2D diffusion models for the colored-PC-to-mesh task, to address the low-quality color reconstruction issue.
+However, it is not trivial to implement this idea: (1) Existing 2D-diffusion-for-3D methods typically focus on 3D generation conditioned on text [15, 16] or images [17, 18], leaving point-cloud-based reconstruction unexplored.
+(2) Existing 2D-diffusion-for-3D methods often yield blurry appearances [19] by predicting colors in 3D space as a texture [20, 16]/radiance [21, 22, 23] field or 3D Gaussians [19, 18].
+How to leverage 2D reference images to generate high-quality textures requires more attention.
+(3) Intermediate 2D multiview images generated by diffusion models often have inconsistencies with each other or with the reconstructed 3D geometry at occlusion border areas.
+When one part of an object occludes another, the generated 2D images and 3D geometry should predict highly consistent occlusion borders.
+Failing to do so would cause artifacts; see “Ours w/o/ NBF” in Figure 1, with significant artifacts near the border of the pillow and the chair.
+We address these challenges one by one.
+First, to bridge the gap between 2D diffusion models and 3D point clouds, our intuition is that, reconstructing 3D textured meshes from colored point clouds is analogous to inpainting 2D sparse images by filling empty pixels.
+They both aim to somehow dreaming the missing areas of the sparse input, to achieve a more complete and coherent representation.
+Luckily, 2D diffusion-based inpainting methods [24] excel even with mask ratios over 80%, making them robust to sparse input points.
+Trained on large-scale 2D datasets, these models generalize well across domains and require no additional training or fine-tuning, enabling zero-shot point cloud reconstruction.
+Second, to address the blurring issue, we propose to directly unproject the generated multiview images onto the 3D mesh, instead of predicting colors in 3D [10, 9] or UV [8] spaces like existing methods.
+Third, to address the occlusion-border inconsistency issue, we design a “Non-Border-First” (NBF) unprojection strategy.
+It detected border areas by leveraging the correspondence of occlusion border in 2D image space and invisibility border in UV space, and then prioritizes non-border views’ images during unprojection, thus avoiding artifacts.
+Putting everything together, we propose PointDreamer, a novel zero-shot framework to reconstruct high-quality textured meshes from colored point clouds, as shown in Figure 1.
+First, we extract geometry (untextured mesh) from the point cloud by an existing method POCO [4].
+Second, we project the point cloud into 2D space from a fixed set of viewpoints, producing multiview sparse images.
+Third, we inpaint the empty pixels with an off-the-shelf 2D diffusion model, forming dense images.
+Finally, we directly unproject the 2D images onto the mesh.
+Unlike most existing methods that predict colors in 3D [10, 11] or UV spaces [8], directly adopting colors of 2D diffusion models’ output high-quality 2D images leads to clear textures.
+In particular, our designed NBF unprojection strategy effectively avoids artifacts caused by inconsistencies in occlusion border regions.
+Overall, we list our contributions below:
+  • We propose PointDreamer, a zero-shot SoTA framework that fills the gap in the less-explored but important task of 3D textured mesh reconstruction from colored point cloud, pioneering in utilizing 2D diffusion models for 3D point cloud reconstruction.
+  • PointDreamer reconstructs high-quality and clear-looking textures for meshes, avoiding the common blurring issue.
+PointDreamer is also zero-shot.
+While requiring no extra training, it shows high generalization capability: robust to noisy, sparse, or even incomplete input points, benefiting from the power of 2D diffusion models.
+  • We design Non-Border-First (NBF) Unprojection, a novel unprojection strategy to apply colors of 2D multiview images onto 3D meshes.
+It effectively avoids artifacts caused by inconsistent prediction near occlusion borders in any multiview-images-to-3D methods.
+Experiments on various benchmarks show the SoTA performance of PointDreamer, by significantly outperforming baseline methods quantitatively and qualitatively.
+2 Related Work
+2.1 Textured Mesh Reconstruction from Colored Point Cloud
+Existing methods can be categorized based on their reliance on training data.
+Non-data-driven methods rely solely on the input point cloud.
+For instance, (Screened) Poisson Surface Reconstruction (PSR, SPSR) [25, 7] reconstructs geometry by solving Poisson equations, and many 3D processing tools [26, 27] extend it with color support via linearly blending input point colors.
+DHSP3D [8] iteratively optimizes a MeshCNN in 3D space and XYZ/RGB maps in UV space by self-supervision.
+On the other hand, data-driven methods train neural networks on 3D datasets to generate textured meshes.
+For example, 3DGen [10] trains a triplane variational autoencoder to predict signed distance field and color values for each 3D tetrahedra grid vertex, and extracts textured meshes by marching tetrahedra [28].
+Hybrid methods employ different strategies for geometry and texture reconstruction.
+NKSR [9] adopts data-driven geometry reconstruction by learning neural kernel fields, while using non-data-driven color reconstruction (in its supplementary file), by optimizing a 3D textured field [12] to overfit the input points’ colors.
+ColorMesh [11] employs non-data-driven geometry reconstruction via graph cuts, paired with a data-driven texture network to inpaint colors in 2D image space.
+However, its texture network is an encoder-decoder CNN trained with limited data.
+This may limit its performance and generalization capability, compared to our adopted zero-shot 2D-diffusion-based inpainting approach trained with extensive 2D data.
+2.2 2D Diffusion Models for 3D Generation
+To capitalize on the abundant availability of 2D data and the high performance of 2D diffusion models, researchers seek to “lift” results of 2D diffusion models to 3D.
+Most of these works optimize or learn a NeRF [22, 23], a DMTet with texture field [29, 16], or 3D Gaussianss by Score Distillation Sampling [13] or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model.
+View-conditioned diffusion models [22] or multi-view diffusion models [30, 14, 31] are developed by fine-tuning existing diffusion models for generating multi-view-consistent images.
+Most of these works focus on 3D generation from text or images instead of point clouds.
+They tend to generate blurry textures, and may suffer from artifacts in occlusion border areas.
+2.3 Border-Area-Inconsistency Issue in 3D Texturing
+Mesh texture reconstruction based on real-scanned or AI-generated multiview images is widely used in applications such as 3D reconstruction and generation [32].
+Due to scanning errors and imperfections in generative models, inconsistencies often arise between different views’ images or between images and the mesh, especially around occlusion borders where one part of the object occludes another.
+However, this issue remains largely unexplored.
+Most texture mapping methods [33, 34, 35, 36] design complex optimization objectives and strategies for camera extrinsics, etc.
+They typically focus on issues like blurring, ghosting, and seams, leaving border-area-inconsistency unexplored.
+Also, they predominantly rely on RGBD data, and thus are unsuitable for scenarios without dense depth data like point cloud reconstruction.
+On the other hand, most 3D generation approaches directly fuse multiview images’ colors for mesh texturing.
+Specifically, they use these images as supervision during optimization [32], or directly compute image colors’ weighted sum as the mesh’s colors [37].
+To address the border-area-inconsistency issue, we design a Non-Border-First unprojection strategy, and it can be used in any method that unprojects multiview images to a mesh.
+3 Method
 Given a 3D point cloud with per-point XYZ coordinates and RGB colors, our goal is to reconstruct its associated textured mesh.
-Unlike existing works that learn or optimize colors in 3D space (Oechsle et al., 2019; Poole et al., 2022) or UV space (Wei et al., 2021; Chen et al., 2022), we propose PointDreamer, a novel approach by conducting color prediction in 2D image space to leverage the powerful 2D diffusion priors.
+To avoid the blurring effect of existing works, we predict colors in 2D space by diffusion-based 2D inpainting, leveraging the powerful 2D diffusion priors.
 Figure 1 presents our designed pipeline with four steps.
-First, we employ a geometry extraction module to reconstruct an untextured mesh from the input point cloud.
-Next, with a set of fixed viewpoints, we perform 3D-to-2D conversion by projecting the input point cloud into 2D, producing sparse multi-view images.
-After that, we conduct color prediction in 2D space by inpainting the empty pixels in the sparse images to form dense ones based on a pre-trained 2D diffusion model (Dhariwal and Nichol, 2021).
-Finally, we propose a novel Non-Border-First strategy to convert the 2D results back to 3D by unprojecting the colors in the dense images to the untextured mesh to produce the desired textured mesh.
-Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training on 3D datasets.
-Figure 2.
-Pipeline of our method.
-Given a colored point cloud, we first reconstruct its untextured mesh.
-Then, the colored point cloud is projected to 2D to produce sparse multi-view images, which are further inpainted to form the dense ones.
-Finally, we unproject the results back to 3D to output the associated textured mesh.
-Figure 3. Illustration of hidden point removal, without which the results would be messed up.
-Figure 4.
-Top: Illustration of input and output of the unprojection module, which unprojects colors of 2D multi-view images to a 3D mesh.
+(1) We employ a geometry extraction module to reconstruct an untextured mesh from the input point cloud.
+(2) With a set of fixed viewpoints, we perform 3D-to-2D conversion by projecting the input point cloud into 2D, producing sparse multi-view images.
+(3) We conduct color prediction in 2D space by inpainting the empty pixels in the sparse images to form dense ones based on a pre-trained 2D diffusion model [38].
+(4) We propose a novel Non-Border-First strategy to convert the 2D results back to 3D by unprojecting the colors in the dense images to the untextured mesh to produce the desired textured mesh.
+Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training.
+It takes only about 61s and 30s for PointDreamer’s DDNM and nearest/linear interpolation variants (see Figure 7 (b)) to reconstruct a shape of 30k points on an NVIDIA A100 GPU, compared to other 2D-diffusion-for-3D methods such as Zero-1-to-3 [22] (~20 m), DreamGaussin [19] (~2 m), Wonder3D [32] (2~3 m), LGM [18] (~65 s), etc.
+Figure 2: Illustration of hidden point removal, without which the results would be messed up.
+Figure 3: Top: Illustration of input and output of the unprojection module, which unprojects colors of 2D multi-view images to a 3D mesh.
 Middle Left: An naive unprojection implementation.
 For each 3D point $\mathbf{p}$, from all views where $\mathbf{p}$ is visible, the highest-direction-priority view is chosen to paint $\mathbf{p}$.
-Example point $\mathbf{p}$’s corresponding pixel in the selected view $a$ is with inaccurate color since it’s near an occlusion border, producing an artifact.
+Example point $\mathbf{p}$’s corresponding pixel in the selected view $a$ is with incorrect color since it’s near an occlusion border, producing an artifact.
 Bottom Right: our proposed Non-Border-First Unprojection.
-Observing the correspondence between occlusion border in 2D image space and visibility in UV space (both marked by sky blue dashed lines), we prioritize non-border-areas in UV space to avoid such artifacts.
-3.1. Geometry Extraction: Point to Surface
+Observing the correspondence between the occlusion border in 2D image space and the visibility border in UV space (both marked by sky-blue dashed lines), we prioritize non-border areas in UV space to avoid such artifacts.
+3.1 Geometry Extraction: Point to Surface
 The first step in PointDreamer is to reconstruct an untextured mesh from the input point cloud.
-Since color information does not need to be considered in this step, many existing point-to-surface methods (Mescheder et al., 2019; Peng et al., 2020; Shen et al., 2021; Boulch and Marlet, 2022) can be employed.
-In our implementation, we directly adopt the state-of-the-art POCO (Boulch and Marlet, 2022) considering its high performance.
-Section 4 provides ablation experiments by using different geometry reconstruction methods.
-3.2. Projection and Inpainting
-As shown in Figure 2, we design a projection module for 3D-to-2D conversion, followed by an inpainting module for 2D color prediction by inpainting sparse multi-view images into dense ones.
-3.2.1. Projection: 3D to 2D
-Directly projecting the whole point cloud into 2D would incorporate points from occluded parts that should not be visible from the given view.
-This would mess up the inpainting results and subsequent mesh textures; see the bottom of Figure 3 as an example.
-Therefore, before generating the sparse images, we precess the input point cloud by the “Hidden Point Removal” operator (Katz et al., 2007), which works by transforming the input and extracting the points that reside on its convex hull.
+Since color information does not need to be considered in this step, many existing point-to-surface methods [39, 40, 28] can be employed.
+In our implementation, we directly adopt the state-of-the-art POCO [^1] https://github.com/valeoai/POCO [4] for its high performance.
+We conduct experiments comparing different geometry reconstruction methods in Section 4 and our supplementary file.
+3.2 Projection: 3D to 2D
+Directly projecting the whole point cloud into 2D space would incorporate points from occluded parts that should not be visible from the given view.
+This would mess up the inpainting results and subsequent mesh textures; see Figure 2 for an example.
+Therefore, before generating the sparse images, we process the input point cloud by the “Hidden Point Removal” operator [41], which works by transforming the input and extracting the points that reside on its convex hull.
+Besides, we compare the depth values of points and the extracted untextured mesh to further remove some invisible points.
 With pre-set camera parameters of $K$ viewpoints ($K=8$ in our implementation) and associated visible point clouds, we conduct camera transformation to these points to get their corresponding pixel coordinates in 2D image space.
 These pixels are painted according to their associated 3D points’ colors, producing $K$ sparse images.
-Comparisons of different $K$ values can be found in our supplementary file.
-3.2.2. Inpainting: Sparse to Dense
+We compare different $K$ values in our supplementary file.
+Algorithm 1 PyTorch-like pseudocode for Non-Border-First (NBF) strategy
+1# K: number of views
+2# T: texture atlas
+3# I: multiview images
+4# I_dep: [K, I_res, I_res], multi-view depth maps obtained from geometry
+5# T_dep: [K, T_res, T_res], T’s pixels’ corresponding depths in each view
+6# Tuv2Iab: [T_res, T_res, 2]: Mapping T’s pixel coordinate to I’s
+7# chart_mask: [T_res, T_res]: Does T(u,v) belong to a chart (in foreground)?
+8# T_normal: [T_res, T_res, 3]: T’s pixels’ corresponding 3D point’ normal directions
+9# view_dir: [K, 3]: camera view directions
+10# I_rgb: [K, I_res, I_res, 3], multi-view RGB images
+11T = torch.zeros(T_res, T_res, 3)
+12vis_T = calc_vis(I_dep, T_dep, Tuv2Iab) # [K, T_res, T_res], bool
+13vis_T_edges = detect_edge(vis_T) # [K, T_res, T_res], bool
+14chart_edges = detect_edge(chart_mask) # [T_res, T_res], bool
+15border_edges = vis_T_edges - chart_edges.unsqueeze(0).repeat(K)
+16border_areas = dilate(border_edges, kernel) # [K, T_res, T_res], bool
+17cand_view_T_map = non_border_areas = ~border_areas # [K, T_res, T_res]
+18Tuv_cand_view_num = non_border_areas.sum(0) # [T_res, T_res], int
+19cand_view_T_map[:, Tuv_cand_view_num==0] = vis_T[:, Tuv_cand_view_num==0]
+20dir_priority = calc_priority(T_normal, view_dir) # [K, T_res, T_res]
+21dir_priority[~cand_view_T_map] -= 1000 # Non-Border-First
+22uv_view_id = dir_priority.argmax(0) # [T_res, T_res]
+23for u in range(0, T_res):
+24 for v in range(0, T_res):
+25 T[u, v] = I_rgb[uv_view_id[u, v], Tuv2Iab[u, v, 0], Tuv2Iab[u, v, 1]]
+3.3 Inpainting: Sparse to Dense
 Any 2D inpainting method that fills empty pixels in input images can serve as our inpainting module.
-We propose to use the state-or-the-art DDNM (Wang et al., 2022b) based on a pre-trained unconditional 2D diffusion model (Dhariwal and Nichol, 2021) to inpaint the multi-view sparse images into dense ones $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$.
-In this way, the strong prior provided by the 2D diffusion model can facilitate high-quality inpainting.
+We propose to use the state-or-the-art DDNM [^2] https://github.com/wyhuai/DDNM  [24] based on a pre-trained unconditional 2D diffusion model [^3] https://github.com/openai/guided-diffusion  [38] to inpaint the multi-view sparse images into dense ones $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$.
+In this way, the 2D diffusion model’s strong prior can facilitate high-quality inpainting.
 So far, we have completed color prediction purely in 2D space instead of 3D or UV space.
-3.3. Unprojection: 2D to 3D
+3.4 Unprojection: 2D to 3D
 With color prediction conducted in 2D space, we now need to convert the result back to 3D.
@@
 The top of Figure 3 illustrates the basic idea of the unprojection module.
-As can be seen, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$ via Xatlas (Young, 2022), following (Gao et al., 2022; Long et al., 2023).
+Specifically, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$.
+We follow existing works [42, 32] to conduct UV mapping by Xatlas [43].
 A texture atlas is an RGB image containing different charts.
@@
 See the top of Figure 3 as an example, where a 3D point $\mathbf{p}$ is visible in two views $a$ and $b$, and we mark it with a red circle together with its corresponding multi-view image pixels $I_{a}[i_{a},j_{a}]$, $I_{b}[i_{b},j_{b}]$, and texture atlas pixel $\mathcal{T}[u,v]$.
-Point $\mathbf{p}$’s visibility in view $k$ can calculated by whether its depth value is no bigger than the corresponding pixel of the $k$’s depth map.
+Point $\mathbf{p}$’s visibility in view $k$ can be calculated by whether its depth value is no bigger than the corresponding pixel of the $k$’s depth map.
 If $\mathbf{p}$ is visible only in one view, the corresponding color of $I_{k}[i_{k},j_{k}]$ can be directly adopted as the color of $\mathbf{p}$ and $\mathcal{T}[u,v]$.
@@
 Various unprojection strategies differ in how to fuse views or select the best view.
-Here we introduce three implementations for unprojection.
-3.3.1. Naive Unprojection: Only Use Direction Priority
+Here we introduce three different implementations for unprojection.
+3.4.1 Naive Unprojection: Only Use Direction Priority
 An intuitive idea to pick the best view for $\mathbf{p}$ or $\mathcal{T}[u,v]$ is by direction priority, as shown in the middle left part of Figure 3.
@@
 While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the middle left part of Figure 3, where a gray artifact is visible in the reconstructed textured mesh that should have been brown.
-Through analysis, such artifacts appear in border areas, that is, the area near the border of occluded (e.g.
-bottom side of the brown seat in Figure 4) and occluding part (e.g.
-the gray bottom of the chair) of the shape in a posed image $I_{k}$ when the former is occluded by the latter from view $k$ (e.g.
-view $a$).
-In such areas, the inpainting module may inaccurately delineate the boundary separating the two parts, e.g.
-too large gray bottom in view $a$, resulting in artifacts.
-3.3.2. Optimization-based Unprojection
-Instead of selecting one single view by direction priority, 3D generation method DreamGaussian (Tang et al., 2023a) adopts an optimization-based unprojection strategy to get the textured mesh from multi-view images.
-The idea is to optimize the texture atlas by the multi-view per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh.
+Through analysis, such artifacts appear in border areas: the area near the border of occluded and occluding part of the shape when viewing from view $k$.
+For example, in view $a$, the brown seat’s bottom side is occluded by the gray bottom of the chair.
+The area near the edge of the brown and gray parts is a border area.
+In such areas, the inpainting module and the geometry extraction module may inconsistently delineate the boundary separating the two parts.
+For example, the gray bottom is larger in the inpainted image $I_{a}$ than of the reconstructed mesh, thus resulting in an artifact.
+3.4.2 Optimization-based Unprojection
+Instead of selecting one single view by direction priority, 3D generation method DreamGaussian [19] adopts an optimization-based strategy to get the textured mesh from multi-view images.
+The idea is to optimize the texture atlas by a per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh.
 In this way, the optimization process kinds of fuses corresponding colors in different views instead of selecting the best one.
-However, it cannot ignore the inaccurate colors in the border areas, so the artifacts can only be reduced instead of eliminated.
+However, it cannot ignore the inconsistent colors in the border areas, so the artifacts can only be reduced instead of eliminated.
 This inspires us to design a non-border-first strategy to address this issue.
-3.3.3. Non-Border-First Unprojection (NBF)
-The bottom of Figure 4 illustrates the idea of our proposed novel unprojection strategy namely Non-Border-First Unprojection, which begins with detecting border areas in multi-view images, so as to paint the texture atlas while ignoring them.
+3.4.3 Non-Border-First Unprojection (NBF)
+The bottom of Figure 3 illustrates the idea of our proposed novel unprojection strategy namely Non-Border-First Unprojection.
+It begins with detecting border areas in multi-view images, so as to paint the texture atlas while deprioritizing them.
 After that, if some atlas pixels remain unpainted since their associated 3D points can not be seen from any non-border area, we then select from all areas, whether they are border areas or not, to paint the rest pixels.
@@
 Therefore, we design our Non-Border-First unprojection strategy as follows.
-  1. (1)
+  1. 1.
     We paint “view $k$ visible texture atlas map” $\mathcal{T}_{k}$ for each view $k$ defined as above.
-  2. (2)
+  2. 2.
     We calculate the border areas by dilating the edges between visible and invisible areas in $\mathcal{T}_{k}$ given a certain dilation kernel size as a hyperparameter.
-  3. (3)
+  3. 3.
     We paint the final texture atlas $\mathcal{T}$ considering only non-border areas in ${\mathcal{T}_{k}}$.
-During this, if a point is visible in more than one view’s non-border-areas, we select the best view with the highest direction priority.
-  4. (4)
-    If there remain some unpainted pixels in $\mathcal{T}$, we now check the previously ignored border areas to paint them, during which we still select among views by direction priority.
-  5. (5)
+During this, if a point is visible in more than one view’s non-border areas, we select the view with the highest direction priority.
+  4. 4.
+    If there remain some unpainted pixels in $\mathcal{T}$, check the previously ignored border areas to paint them, during which we still select among views by direction priority.
+  5. 5.
     If some points cannot be seen from either view, we assign it with a view by direction priority considering all areas in all views whether they’re border areas or not.
-  6. (6)
-    We experimentally find that additionally optimizing from the generated texture atlas by only calculating the loss of non-border pixels would further enhance the performance.
-In this way, our non-border-first strategy can reasonably eliminate artifacts caused by inaccurate color prediction in border areas.
+  6. 6.
+    We experimentally find that additionally optimizing from the generated texture atlas by only calculating the loss of non-border pixels would further enhance the results.
+In this way, our non-border-first strategy can reasonably eliminate artifacts caused by inconsistent prediction in border areas.
+Altorithm 1 shows the corresponding pseudoCode.
 We conduct experiments to verify the effectiveness of our proposed strategy in Section 4.
-4. Experiments and Results
-4.1. Experimental Settings
-4.1.1. Datasets
-To evaluate the performance of our method against other competitors, we use three benchmark datasets in our experiments.
-(i) ShapeNetCoreV2 (Chang et al., 2015): A large-scale synthesis 3D object dataset, where we follow (Gao et al., 2022) to use the official test splits of chair, car, and motorbike categories for evaluation since they contain relatively complex textures.
-The test sets of the three categories contain around 1300, 690 and 70 samples, respectively.
-(ii) Google Scanned Objects (GSO) (Downs et al., 2022): A real-scanned 3D object dataset and we use all its 1030 samples for evaluation unlike existing works (Liu et al., 2024; Long et al., 2023) that only use 30 of them.
-(iii) OmniObject3D (Wu et al., 2023): A real-scanned 3D object dataset with 6000 samples. For efficiency, we randomly select 100 objects for evaluation.
-For each textured mesh from the above datasets, we sample 30000 colored points as input.
-Note that, our approach works in a zero-shot manner, which can be directly applied to objects from various datasets or categories, requiring no extra training but only an off-the-shelf 2D diffusion model (Dhariwal and Nichol, 2021).
-4.1.2. Metrics
-To evaluate the quality of the reconstructed 3D textured mesh, we follow existing works (Liu et al., 2023c, 2024; Long et al., 2023; Zhang et al., 2022) to compare the similarity between multi-view images rendered by the reconstructed mesh and ground-truth mesh.
-Specifically, we render 20 multi-view images on each mesh for a thorough assessment and four similarity metrics are calculated: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) (Wang et al., 2004), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), and Fréchet Inception Distance (FID)  (Heusel et al., 2017).
-PSNR measures pixel-level similarity.
-SSIM considers both pixel values and structural information.
-LPIPS uses a pre-trained deep neural network to assess image similarity, imitating human perception better than PSNR and SSIM.
-FID measures the statistical distance between the distribution of two datasets.
-PSNR and SSIM scores indicate higher quality as their values increase, whereas lower values of LPIPS and FID reflect better performance.
-4.2. Baselines
-To fully evaluate the performance of our method, we compare it with three kinds of baselines:
-(1) Non-deep-learning method: We choose the classical Screened Poisson Reconstruction (SPR) (Kazhdan and Hoppe, 2013) for comparison.
-Note that SPR requires per-point normal as input, and we estimate it via Principal Component Analysis (PCA).
-(2) DHSP3D (Wei et al., 2021) and NKSR (Huang et al., 2023): These are two recent networks for textured mesh reconstruction and we employ the official code provided by the authors.
-Similar to SPR, NKSR also requires point normals as additional input.
-(3) Texture Field (Oechsle et al., 2019): Since we cannot find any existing open-source texture-field-based method requiring only colored point clouds as input, we thus design a baseline network inspired by (Gupta et al., 2023).
+4 Experiments and Results
+4.1 Experimental Settings
+4.1.1 Datasets
+To evaluate the performance of our method against other competitors, we use three benchmark datasets:
+(i) ShapeNetCoreV2 [44]: A large-scale synthesis 3D object dataset, where we follow [42] to use the official test splits of chair, car, and motorbike categories for evaluation, since they contain relatively complex textures.
+The test sets of the three categories contain around 1300, 690, and 70 samples, respectively.
+(ii) Google Scanned Objects (GSO) [45]: A real-scanned 3D object dataset and we use all its 1030 samples for evaluation, unlike existing works [14, 32] that only use 30 of them.
+(iii) OmniObject3D [46]: A real-scanned 3D object dataset with 6000 samples. For efficiency, we randomly select 100 objects for evaluation.
+Figure 4: Qualitative results on (a) ShapeNetCoreV2 dataset (Chair, Car, Motorbike Category), (b) GSO dataset, (c) OmniObject3D dataset. (d) Comparison with DHSP on GSO. (e) Our Multiview Consistency.
+For each textured mesh from the above datasets, we sample 30k colored points as input.
+Note that, our approach is zero-shot and can be directly applied to objects from various datasets or categories, requiring no extra training but only an off-the-shelf 2D diffusion model [38].
+4.1.2 Metrics
+To evaluate the quality of the reconstructed 3D textured mesh, we follow existing works [22, 14, 32, 47] to compare the similarity between multi-view images rendered by the reconstructed mesh and ground-truth mesh.
+Specifically, we render 20 images per shape for a thorough assessment by distributing cameras on the vertices of a regular icosahedron.
+Four similarity metrics are calculated: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [48], Learned Perceptual Image Patch Similarity (LPIPS) [49], and Fréchet Inception Distance (FID)  [50].
+PSNR and SSIM focus more on pixel-level similarity, while LPIPS and FID imitate human perception better.
+Since we focus on texture instead of geometry reconstruction, we put geometry evaluation results in our supplementary file.
+4.2 Baselines
+To fully evaluate the performance of our method, we compare it with three kinds of baseline methods:
+(1) Screened Poisson Reconstruction (SPR) [7]: SPR is a classical mesh reconstruction method.
+We adopt the implementation of the commonly-used tool MeshLab [26], which generates texture by linearly blending colors of the input point cloud.
+Note that SPR requires per-point normal as input, which we estimate via MeshLab’s default Principal Component Analysis (PCA).
+(2) DHSP3D [8] and NKSR [9]: These are two recent approaches for mesh reconstruction which generate textures by overfitting the input points’ colors.
+We employ the official code provided by the authors  [^4] https://github.com/weixk2015/DHSP3D,
+https://github.com/nv-tlabs/NKSR/blob/public/examples/
+recons_colored_mesh.py
+. Similar to SPR, NKSR also requires point normals as additional input.
+(3) Texture Field (T.F.)  [12]: Since we cannot find any existing open-source texture-field-based method requiring only colored point clouds as input, we thus design a baseline network inspired by [10].
 The key idea is to learn a 3D feature tri-plane from the input point cloud and then decode RGB values for the query points.
-We train it on the official training split of the mentioned three categories of ShapeNetCoreV2, supervised by per-point color MSE loss following (Gupta et al., 2023).
-During inference, we employ the above-mentioned POCO (Boulch and Marlet, 2022) to reconstruct untextured meshes and utilize our trained texture field network to infer textures by querying colors for the associated 3D point of each texture atlas pixel.
-In this way, both our proposed PointDreamer and the texture field baseline adopt the same geometry module for a fair comparison.
-Please refer to our supplementary file for more details about the training and inference of our texture field baseline.
-Figure 5. Qualitative results of textured mesh reconstruction on ShapeNetCoreV2 dataset.
-Figure 6. Qualitative results of textured mesh reconstruction on Google Scanned Objects dataset.
-Figure 7. Qualitative results of textured mesh reconstruction on OmniObject3D dataset.
-[…]
-Figure 8. Qualitative results on textured mesh reconstruction comparing our method and DHSP3D.
-Table 1. Comparisons on textured mesh reconstruction performance on ShapeNetCoreV2 dataset.
+We train it on the official training split of the mentioned three categories of ShapeNetCoreV2, supervised by per-point color MSE loss following [10].
+During inference, we employ the above-mentioned POCO [4] to reconstruct untextured meshes.
+Then we utilize our trained T.F.
+network to infer textures, by querying colors for the associated 3D point of each texture atlas pixel.
+In this way, both PointDreamer and the T.F.
+baseline adopt the same geometry module for a fair comparison.
+Please refer to our supplementary file for more details about the training and inference of our T.F.
+baseline.
+TABLE I: Comparisons on textured mesh reconstruction on ShapeNetCoreV2, GSO and Omniobject3D datasets.
+T.F.
+refers to Texture Field.
+Note that both T.F.
+and our PointDreamer adopt POCO [4] for geometry reconstruction for a fair comparison.
 {NiceTabular}
 c—c—cccc ShapeNet Cat. Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
-Chairs SPR 24.10 0.931 0.0923 12.03
-NKSR 23.05 0.931 0.0964 29.11
-T.F. 26.12 0.947 0.0664 6.83
-Ours 26.29 0.952 0.0570 4.90
-Cars SPR
-23.46 0.929 0.0882 43.89
-NKSR 21.01 0.919 0.0961 131.05
-T.F. 22.42 0.919 0.0868 38.08
-Ours 22.78
-0.930 0.0727 12.41
-Motorbikes SPR 15.09 0.809 0.2139 106.09
-NKSR 18.13 0.905 0.1008 175.42
-T.F. 21.17 0.926 0.0635 40.06
-Ours
-21.20 0.930 0.0562 28.42
-Table 2. Comparisons on textured mesh reconstruction performance on GSO and Omniobject3D datasets. “T.F” denotes Texture Field.
-{NiceTabular}
-c—c—cccc Dataset Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
-GSO SPR 27.11 0.907 0.1200 27.43
-NKSR 24.65 0.900 0.1184 36.51
-T.F. 25.53 0.894 0.1276 46.44
-Ours 27.24 0.923 0.0827 9.32
-OmniObject3D SPR
-30.20 0.927 0.0998 40.13
-NKSR 26.77 0.919 0.1048 50.14
-T.F. 28.08 0.914 0.1121 65.61
-Ours 29.78
-0.941 0.0683 18.23
-4.3. Main Results
-4.3.1. Qualitative comparisons
-Since DHSP3D is much slower than other methods (about 4-6 hours per shape compared to ¡ 2 min), we randomly select two objects from GSO dataset for efficiency when comparing our method against DHSP3D; see Figure 8.
-For other methods (i.e., SPR, NKSR & texture field), we present the visual comparisons on each dataset separately; see Figures 56 and 7.
-Figures 57 show that our method, though zero-shot, achieves much clearer and more realistic textures than baselines, thanks to the strong 2D diffusion prior.
+Chairs SPR 24.10 0.931 0.092 12.03
+NKSR 23.05 0.931 0.096 29.11
+T.F. 26.12 0.947 0.066 6.83
+Ours 26.29 0.952 0.057 4.90
+Cars SPR 23.46 0.929 0.088 43.89
+NKSR 21.01 0.919 0.096 131.05
+T.F. 22.42 0.919 0.087 38.08
+Ours 22.78 0.930 0.073 12.41
+Motorbikes SPR 15.09 0.809 0.214 106.09
+NKSR 18.13 0.905 0.101 175.42
+T.F. 21.17 0.926 0.064 40.06
+Ours 21.20 0.930 0.056 28.42
+Dataset Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
+GSO SPR 27.11 0.907 0.120 27.43
+NKSR 24.65 0.900 0.118 36.51
+T.F. 25.53 0.894 0.128 46.44
+Ours 27.24 0.923 0.083 9.32
+OmniObject3D SPR 30.20 0.927 0.100 40.13
+NKSR 26.77 0.919 0.105 50.14
+T.F. 28.08 0.914 0.112 65.61
+Ours 29.78 0.941 0.068 18.23
+4.3 Main Results
+4.3.1 Qualitative Results
+We present visual results in Figure 4.
+Since DHSP3D is much slower than other methods (4-6 hours per shape compared to $<$1 min), we randomly select two objects from GSO dataset for efficiency during comparison.
+We highly recommend referring to our supplementary materials for more visual results and a 360° video.
+Figure 4 (a)-(d) show that our method, though zero-shot, achieves much clearer and more realistic textures than baselines, thanks to the 2D diffusion prior.
 Besides, SPR and NKSR, which both rely on per-point normals as input, sometimes yield redundant geometry mainly due to the wrongly-estimated normal direction.
 The texture field baseline, which is trained on ShapeNetCoreV2, performs much worse on the newly-seen dataset of GSO and Omniobject3D, indicating a lack of generalization ability.
-In Figure 8, the red cup in the first row shows that DHSP3D does not support arbitrary topology.
+In Figure 4 (d), the red cup in the first row shows that DHSP3D does not support arbitrary topology.
 This is because it optimizes meshCNN and 2D XYZ map from the convex hull of the input point cloud, enforcing the output mesh to adhere to the hull’s topology.
 Besides, DHSP3D suffers from less realistic textures with jagged or unclear edges, since its purely self-prior property only enables it to predict point colors considering corresponding neighbors, without the ability to dream the unseen like our PointDreamer.
-4.3.2. Quantitative comparisons
-We summarize the quantitative results in Tables 4.2 and 4.2.
-We can see that our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin.
+4.3.2 Quantitative comparisons
+We summarize the quantitative results in Tables 4.2, and geometry evaluation results in our supplementary file.
+Our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin.
 However, regarding PSNR, which prioritizes pixel-level accuracy and thus may differ from human perception of quality, our method is sometimes outperformed by SPR.
-We assume that SPR predicts a point’s color by somehow fusing nearby points’ colors, leading to blurry textures but safely avoiding extreme pixel-level errors.
+PSNR has been found to prefer blurry images [49].
+We assume that SPR predicts a point’s color by blending nearby points’ colors, leading to blurry textures but avoiding extreme pixel-level errors.
 On the contrary, our method dreams the colors by its diffusion prior.
-While achieving visually better results, a few pixels with extremely high errors may lower the overall PSNR score.
-4.4. Impact of Degraded Input Quality
-4.4.1. Anti-Noise Ability Analysis
-We compare the anti-noise abilities of our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset.
-Table 4.5 shows the results, where our method, even with noisy input, outperforms baselines with clean input, indicating a strong anti-noise ability.
-Visual comparisons can be found in our supplementary file.
-4.4.2. Sparsity Test
-To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 20k gradually, and present the results in Table 4.5.
-We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input; see results in Table 4.2 “Chairs” category.
-Again, visual comparisons can be found in our supplementary file.
-4.5. Method Component Analysis
-We here compare different implementations for sub-modules (geometry extraction, inpainting and unprojection module) of our method by replacing one module at a time while keeping others unchanged.
+While achieving visually better results, a few pixels with extremely high errors may lower the PSNR score.
+4.4 Multi-view Consistency
+Figure 4 (e) and our supplementary video show that, our method achieves high multi-view consistency.
+The reason is two-fold: (1) Unlike image-conditioned 3D generation, it’s easier to maintain multi-view consistency when reconstructing point clouds, since the input point cloud already determines the global shape and colors, leaving only local details to be dreamed by the model.
+(2) For inconsistencies in local details, our proposed Non-Border-First unprojection effectively addresses this issue; see Figure 7 (c) and Table 4.4.
+TABLE II: Anti-noise comparisons on ShapeNetCoreV2 Chairs. ‘Noisy’ means adding Gaussian noise with a standard deviation of 0.005 to the input.
+{NiceTabular}
+cc—cccc
+Input Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
+Noisy SPR 19.64 0.884 0.170 65.25
+Noisy NKSR 22.79 0.929 0.102 44.02
+Noisy T.F. 26.01 0.944 0.071 8.21
+Noisy Ours 26.26 0.952 0.057 4.93
+Clean T.F. 26.12 0.947 0.066 6.83
+TABLE III: Performance comparison on ShapeNetCoreV2 Chairs with increased input point cloud sparsity.
+{NiceTabular}
+cc—cccc
+Point Number Method PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
+30k Ours 26.29 0.9524 0.057 4.90
+25k Ours 26.28 0.9517 0.058 5.24
+20k Ours 26.26 0.9509 0.059 5.70
+10k Ours 26.17 0.9477 0.064 7.79
+30k T.F. 26.12 0.9467 0.066 6.83
+TABLE IV: Quantitative results of replacing our NBF unprojection strategy with others. “Opt.” denotes “Optimize”.
+{NiceTabular}
+c—cccc Unprojection Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
+Opt. Scratch 26.19 0.9473 0.0592 5.26
+Naive 26.22 0.9500 0.0587 5.11
+Opt. Naive 26.25 0.9511 0.0580 4.94
+NBF 26.27 0.9512 0.0579 5.03
+Opt. NBF 26.26 0.9516 0.0574 4.93
+Figure 5: Our PointDreamer’s robustness against noisy and sparse input.
+Figure 6: Our PointDreamer’s unique completion ability.
+When the input point cloud is incomplete e.g.
+scanned from occluded objects as shown in (a), only our PointDreamer can plausibly complete the missing part with high-quality texture, as demonstrated in (c).
+This is benefited from the strong inpainting power of the adopted 2D diffusion model, as illustrated in (b).
+Figure 7: Sub-Module Replacement Results: (a) Geometry module. (b) Inpainting module. (c) Unprojection module.
+Figure 8: Examples of adopting our proposed NBF strategy on other methods (InstantMesh [51]).
+(a) The original mesh generated by InstantMesh looks blurry.
+(b) Directly unprojecting intermediate multiview images to the mesh alleviates the blurriness, but leads to artifacts at the border of the eye as marked in a red box.
+(c) Using our NBF unprojection strategy avoids such artifacts.
+(d-g) More examples.
+4.5 Robustness against Degraded Input Quality
+4.5.1 Anti-Noise Ability Analysis
+We compare our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset.
+Table 4.4 shows the results, where our method, even with noisy input, outperforms baselines with clean input.
+Figure 5 (b) also shows our PointDreamer’s strong anti-noise ability.
+While baseline methods exhibit significant performance drops given noisy input, our PointDreamer’s performance degradation is minimal: even barely noticeable by human perception.
+4.5.2 Sparsity Test
+To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 10k gradually, and present the results in Table 4.4.
+We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input.
+When the input point number is decreased to 10k, our PointDreamer achieves higher PSNR and SSIM metrics compared to the Texture Field baseline with 30k points.
+We also provide visual comparisons in Figure 5 (c), where our method significantly outperforms baseline methods with only 10k points as input.
+4.5.3 Texture Completion Ability
+Thanks to the inpainting power of the adopted 2D diffusion model, our PointDreamer has a unique ability to complete textures from incomplete scans.
+Figure 6 shows an example, where the point cloud of the table is incomplete since a small area is occluded by the cup during scanning.
+As shown in Figure 6 (a), the tablecloth is partially missing, revealing the black underside of the table below.
+For geometry reconstruction, such incompletion is less challenging: both our method and baseline methods manage to fill in this hole, as shown in Figure 6 (c).
+If facing more challenging cases, we can also add a point cloud completion module [52] before mesh reconstruction.
+However, texture completion is less explored, and baseline methods cannot effectively deal with it.
+Luckily, as shown in Figure 6 (b), our diffusion-based inpainting module addresses this issue.
+When projecting the incomplete 3D point cloud into a 2D sparse image from the top view, the occluded part is empty to be inpainted, thanks to our hidden point removal operation.
+The strong diffusion prior handles this empty area well and the inpainted dense image looks plausible.
+As a result, as shown in Figure 6 (c), our PointDreamer completes the missing area with high-quality texture compared to other methods.
+4.6 Method Component Analysis
+We here compare different implementations for sub-modules (geometry extraction, inpainting, and unprojection module) of our method by replacing one module at a time while keeping others unchanged.
 All experiments in this subsection are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
-Table 3. Comparisons on anti-noise ability on chair category of ShapeNetCoreV2 dataset. “Noisy” denotes adding Gaussian noise with a standard deviation of 0.005 to the input point cloud.
-{NiceTabular}
-c—cc—cc—cc—cc
-Method PSNR ↑   SSIM ↑   LPIPS ↓   FID ↓
-Clean Noisy Clean Noisy Clean Noisy Clean Noisy
-SPR 24.10 19.64 0.9311 0.8840 0.0923 0.1702 12.03 65.25
-NKSR 23.05 22.79 0.9307 0.9288 0.0964 0.1017 29.11 44.02
-T.F. 26.12 26.01 0.9467 0.9436 0.0664 0.0711 6.83 8.21
-Ours 26.29 26.26 0.9524 0.9516 0.0570 0.0574 4.90 4.93
-Table 4. Comparisons on our method’s performance with increased input point cloud sparsity.
-{NiceTabular}
-c—cccc Point Number PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
-30k 26.2910 0.9524 0.0570 4.9048
-25k 26.2761 0.9517 0.0580 5.2412
-20k 26.2553 0.9509 0.0594 5.6960
-4.5.1. Geometry Extraction
-In addition to POCO, we further explore extracting geometry by SPR and depth inpainting to investigate how the geometry extraction module influence the final reconstructed textured mesh.
-Depth inpainting refers to inpainting projected sparse 2D depth maps instead of RGB images, and then reconstructing geometry from the inpainted dense depth maps by depth fusion (Griegler, 2017).
-Please refer to our supplementary file for more details and the quantitative results.
-We present the visual comparisons in Figure 9.
-We can see that, compared to POCO, both SPR and depth inpainting suffer from noisy geometry, resulting in noisy textures.
-Clearly, a higher-performance geometry extraction module contributes to a more refined reconstructed textured mesh.
-Figure 9. Geometry extraction module replacement results comparing our POCO against SPR and Depth Inpainting.
-4.5.2. 2D Inpainting
-We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR (Zhu et al., 2023) and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability.
-We provide the visual comparisons in Figure 10 and quantitative results in our supplementary file.
+4.6.1 Geometry Extraction
+In addition to POCO, we further explore extracting geometry by SPR and depth inpainting to investigate how geometry quality influences the final reconstructed textured mesh.
+Please refer to our supplementary file for more details on depth inpainting and quantitative results.
+Figure 7 (a) shows that noisy geometry leads to noisy texture, and our adopted POCO performs the best.
+4.6.2 2D Inpainting
+We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR [53], and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability.
+We provide the visual comparisons in Figure 7 (b) and quantitative results in our supplementary file.
 As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, indicating the importance of a strong inpainting module to the reconstruction performance.
-Figure 10. Inpainting module replacement results comparing our adopted DDNM against nearest interpolation, linear interpolation, and DiffPIR.
-4.5.3. Unprojection
-We present the comparison results of different unprojection strategies in Figure  11 and Table 4.5.3.
-  • “Opt. Scratch” denotes optimizing the texture atlas from scratch (Tang et al., 2023a);
-  • “Naive” selects the best views by direction priority;
-  • “NBF” selects the best views by our proposed Non-Border-First strategy;
-  • “Opt. Naive” further optimizes the obtained atlas of “Naive”;
-  • “Opt. NBF” further optimizes the obtained atlas of “NBF” while considering only loss in non-border areas.
-The first row of Figure 11 shows an example, where optimizing from scratch fails on areas unseen from any view due to occlusion.
-Both rows show that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization.
+4.6.3 Unprojection
+We present the comparison results of different unprojection strategies in Figure  7 (c) and Table 4.4.
+  • Opt. Scratch: optimize the texture atlas from scratch;
+  • Naive: select the best views by direction priority;
+  • NBF: select the best views by our proposed Non-Border-First strategy;
+  • Opt. Naive: further optimize Naive’s obtained atlas;
+  • Opt. NBF: further optimizes NBF’s obtained atlas while considering only loss in non-border areas.
+Figure 7 (c) shows that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization.
 Only our NBF and Opt.
 NBF nearly eliminate such artifacts.
-Table 4.5.3 also shows that our proposed NBF and Opt.
+Table 4.4 also shows that our proposed NBF and Opt.
 NBF outperform other unprojection strategies.
-Figure 11. Qualitative results of replacing our unprojection strategy with others. “Opt.” denotes “Optimize”.
-Table 5. Quantitative results of replacing our NBF unprojection strategy with others. “Opt.” denotes “Optimize”.
-{NiceTabular}
-c—cccc Unprojection Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
-Opt. Scratch (Tang et al., 2023a) 26.1939 0.9473 0.0592 5.2619
-Naive 26.2225 0.9500 0.0587 5.1068
-Opt. Naive 26.2477 0.9511 0.0580 4.9389
-NBF 26.2729 0.9512 0.0579 5.0309
-Opt. NBF 26.2565
-0.9516 0.0574 4.9326
-5. Conclusion, Limitations and Future Work
+Our NBF strategy can also work for other methods [51, 37] that reconstruct mesh textures from multiview images.
+For example, InstantMesh [51] reconstructs a textured 3D mesh from a single image, by first generating multiview images, and then predicting the mesh by a large reconstruction model.
+As mentioned, predicting colors by a network trained with 3D data can result in blurry textures, as shown in Figure 8 (a).
+While unprojecting the multiview images to the mesh can achieve clearer textures, as shown in Figure 8 (b), artifacts often occur at border areas, such as the border of the eye as marked in red.
+Our NBF strategy effectively solved such artifacts, as shown in (c).
+More examples are provided in (d-g).
+5 Conclusion, Limitations and Future Work
 Conclusion.
-We propose PointDreamer, a novel zero-shot framework for textured mesh reconstruction from colored point cloud.
-PointDreamer predicts colors by 2D inpainting to utilize the strong diffusion prior and achieves SoTA performance.
+We propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud with SoTA performance.
+By utilizing diffusion-based 2D inpainting, it (1) reconstructs clear and high-quality mesh textures, addressing the common blurring issue; (2) shows high robustness against sparse, noisy, or even incomplete input, and (3) works in a zero-shot manner, requiring no extra training.
 We also propose a novel “Non-Border-First” strategy to unproject the colors of predicted 2D images back to 3D space.
-Limitation and Future Work.
-The main shortcoming of our proposed method is its relatively low speed.
-We employed one NVIDIA GeForce GTX 3090 GPU and it takes about 100s for our method to reconstruct a textured mesh, compared to NKSR’s 0.64 s.
-It takes about 72 s for DDNM (Wang et al., 2022b) to inpaint our 8 multi-view images, which takes most of the time.
-Replacing DDNM with a faster inpainting module may increase our reconstruction speed.
-In the future, we may seek to address this issue, and further explore 3D-aware inpainting and relighting-supported texture reconstruction.
-Appendix A Appendix: Implementation Details
-A.1. Texture Field Baseline
-A.1.1. Network architecture
-Inspired by works (Gupta et al., 2023; Shen et al., 2021) that represent 3D information by a feature tri-plane, we adopt the network architecture of the open-source Convolutional Occupancy Network (Peng et al., 2020), which also follows a tri-plane representation.
-Specifically, we modify its one-dimensional output head for occupancy prediction to three-dimensional for RGB color prediction.
-A.1.2. Training
-We follow 3DGen (Gupta et al., 2023) to train the network by the per-3D-point MSE loss on predicted and GT colors of sampled 3D points.
-Additionally, we also tried to employ the per-2D-pixel MSE loss of 2D images rendered from GT meshes and predicted meshes by differentiable rendering, but the experimental results are worse, as shown in Figure 12.
-Therefore, we adopt the per-3D-point MSE loss in our manuscript.
-We train the texture field network on a single NVIDIA GeForce GTX 3090 GPU with a batch size of 24 and a learning rate of 0.0001 for 2592, 024 iterations (about 356 epochs).
-A.1.3. Inference.
-During inference, we first use POCO (Boulch and Marlet, 2022) to predict an untextured mesh from the input point cloud, and apply UV mapping to it by Xatlas (Young, 2022), which produces the 3D positions of each valid pixel in the texture atlas.
-We query the color of each of these 3D positions by our trained texture field network, to inpaint the texture atlas, so as to obtain the final textured mesh.
-Figure 12.
-Visual comparisons of the Texture Field baseline trained with different losses.
-“2D per-pixel” denotes rendering the generated textured mesh to multi-view 2D images, and then calculating the MSE loss between the rendered and GT images.
-“3D per-point” denotes calculating MSE loss between the predicted and GT colors of sampled 3D points.
-A.2. Depth Inpainting for Geometry Extraction
+This strategy is the first to address the border-area artifact issue, which is less-explored but commonly-occurred in methods that generate 3D textures from multiview images.
+Limitations and Future Work.
+(1) Like all other methods producing meshes from multi-view images, our method cannot perfectly color the small unseen areas.
+(2) The adopted inpainting module DDNM, though with SoTA performance, cannot perfectly inpaint all cases, especially those with extremely complex, irregular patterns or very fine details.
+In the future, we plan to develop an adaptive camera placement strategy to minimize invisible regions and better cover the entire mesh.
+Additionally, we aim to further investigate a relighting-supportive reconstruction approach: disentangling illumination effects from the input points’ colors, and reconstructing albedo colors and material information for the mesh, to better support relighting in graphics pipelines.
+Appendix A Geometry Evaluation Results
+To compare the reconstructed geometry quality of PointDreamer and baseline methods while ignoring textures, we report geometry evaluation results in Table F.
+Same as our manuscript, we use ShapeNetCoreV2 [44] (Chair, Car, Motorbike categories), Google Scanned Objects [45] and Omniobject3d [46] datasets.
+We follow POCO [4] to use commonly-used metrics of Chamfer L1-distance ×100 (CD), normal consistency (NC), and F-Score with threshold value 1% (FS).
+See POCO for more details about these metrics.
+Note that, as mentioned in our manuscript, we use POCO as the geometry extraction module for both our PointDreamer and the Texture Field baseline.
+Appendix B More Visual Results
+Figure 10 shows more visual comparisons of our PointDreamer against baseline methods.
+Figure 11 shows more visual results of our PointDreamer, including projected sparse multi-view images and inpainted dense multi-view images.
+We hope this can help the readers better understand our pipeline.
+Appendix C Method Component Analysis
+C.1 Geometry Extraction Module Replacement Experiment
+To investigate how the geometry extraction module influences the final reconstructed textured mesh, we explore three different geometry extraction methods: POCO [4] (as adopted in our manuscript), SPR [7] and depth inpainting.
+C.1.1 Depth-Inpainting for geometry extraction
 With 2D inpainting adopted as the key for our texture reconstruction, we can also introduce it to geometry reconstruction, by inpainting 2D depth maps instead of RGB images.
-Specifically, we follow the following steps to reconstruct untextured meshes from the input point cloud, as shown in Figure 13:
-Figure 13. Pipeline of extracting geometry by depth inpainting.
-  1. (1)
+Regarding this, we conduct experiments of depth inpainting, which refers to inpainting projected sparse 2D depth maps instead of RGB images, and then reconstructing geometry from the inpainted dense depth maps by depth fusion [54].
+Specifically, we follow the following steps to reconstruct untextured meshes from input point clouds, see Figure 12:
+  1. 1.
     We generate multi-view sparse depth maps by projecting 3D points to 2D, and assigning the value of each pixel as the depth value of the corresponding 3D point.
 Note that, similar to generating our sparse RGB images, hidden point removal is conducted for each viewpoint before projecting.
-  2. (2)
+  2. 2.
     Since depth fusion requires depth maps’ background pixels to have infinite values to produce a reasonable mesh, we generate foreground masks by projecting 3D points to 2D space with a relatively big point size, i.e.
@@
 In addition, since we use a big point size to generate the foreground mask, the mask would be bigger than the ground truth, so we shrink the generated mask by erosion.
-  3. (3)
+  3. 3.
     We inpaint the foreground pixels of the sparse depth maps into dense ones by nearest interpolation considering efficiency.
-  4. (4)
-    We produce an untextured mesh by depth fusion based on the inpainted dense depth maps, and conduct mesh simplification (Garland and Heckbert, 1997) and Taubin Smooth (Taubin, 1995) to it as post-processing, to get the final untextured mesh.
-Appendix B Appendix: More Experimental Results
-B.1. Effect of Different $K$ Values (Number of Viewpoints for Projection and Inpainting)
+  4. 4.
+    We produce an untextured mesh by depth fusion based on the inpainted dense depth maps, and conduct mesh simplification [55] and Taubin Smooth [56] to it as post-processing, to get the final untextured mesh.
+C.1.2 Results
+We present the visual comparisons in our manuscript and quantitative results in Table F, respectively.
+As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID.
+Both SPR and Depth Inpainting suffer from noisy geometry and thus low-quality textures.
+This indicates that a higher-performance geometry extraction module contributes to a more refined reconstructed textured mesh.
+C.2 2D Inpainting Module Replacement Experiment
+We provide the quantitative results of replacing our 2D inpainting module from DDNM [24] to nearest interpolation, linear interpolation, and DiffPIR [53] in Table F.
+DiffPIR is another 2D-diffusion-based image restoration approach with inpainting ability.
+As can be seen, DDNM outperforms the other three methods thanks to its strong inpainting ability, which is consistent with the visual comparisons from Figure 4 (b) in our manuscript.
+Appendix D More Experimental Results
+D.1 Effect of Different K Values (Number of Viewpoints for Projection and Inpainting)
 To investigate the effect of different numbers of viewpoints (denoted as $K$) for projection and inpainting, we conduct experiments on the motorbike category or ShapeNetCoreV2 dataset by using different $K$ values including 6, 8, and 20.
-Table B.1 shows the distribution of cameras for each setting, together with the quantitative results.
+Table F shows the distribution of cameras for each setting, together with the quantitative results.
 We can see that more views contribute to a slightly higher reconstruction quality.
@@
 The only exception is the motorbike category from the ShapeNetCoreV2 dataset, for which we use the 20 views instead, considering motorbikes’ more complex geometry and topology.
-Table 6.
-Quantitative results of using different $K$ values (numbers of viewpoints for projection and inpainting) on the motorbike category of ShapeNetCoreV2 dataset.
+D.2 Impact of Degraded Input Quality: More Visual Comparisons
+D.2.1 Anti-Noise Ability Analysis
+Figure 14 presents the visual comparisons of our method and baseline methods’ reconstructed meshes from noisy or clean input point clouds.
+Note that the reconstructed untextured meshes of noisy and clean inputs are the same, this is because our adopted POCO was trained with noisy input, so before reconstructing geometry, we manually add noise to the clean input.
+Overall, we have the following observations, which are consistent with the quantitative results (Table 2) in our manuscript:
+  1. 1.
+    As expected, all methods produce higher-quality textured meshes with clean input point clouds compared to noisy ones.
+  2. 2.
+    Our PointDreamer shows a relatively high anti-noise ability, where only a small performance drop is observed when giving noisy input.
+  3. 3.
+    Our PointDreamer with noisy input point clouds shows an even better visual effect compared to baseline methods with clean inputs.
+D.2.2 Sparsity Test
+Figure 15 shows the visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input.
+There is a very small performance drop introduced by decreasing the input point number, which can sometimes be hard to notice by human eyes.
+This indicates a relatively high robustness of our method towards varying degrees of input sparsity.
+Appendix E Implementation Details of Texture Field Baseline
+E.1 Network architecture
+Inspired by works [10, 28] that represent 3D information by a feature tri-plane, we adopt the network architecture of the open-source Convolutional Occupancy Network [40], which also follows a tri-plane representation.
+Specifically, we modify its encoder from taking three-dimensional input (xyz) to six-dimensional input (xyzrgb).
+Then, we modify its one-dimensional output head for occupancy prediction to three-dimensional for RGB color prediction.
+E.2 Training
+We follow 3DGen [10] to train the network by the per-3D-point MSE loss on predicted and GT colors of sampled 3D points.
+Additionally, we also tried to employ the per-2D-pixel MSE loss of 2D images rendered from GT meshes and predicted meshes by differentiable rendering, but the experimental results are worse, as shown in Figure 16.
+Therefore, we adopt the per-3D-point MSE loss in our manuscript.
+We train the texture field network on a single NVIDIA GeForce GTX 3090 GPU with a batch size of 24 and a learning rate of 0.0001 for 2592, 024 iterations (about 356 epochs).
+E.3 Inference.
+During inference, we first use POCO [4] to predict an untextured mesh from the input point cloud, and apply UV mapping to it by Xatlas [43], which produces the 3D positions of each valid pixel in the texture atlas.
+We query the color of each of these 3D positions by our trained texture field network, to inpaint the texture atlas, so as to obtain the final textured mesh.
+Appendix F Applications for large scenes
+PointDreamer, as a zero-shot method, can be adapted to indoor or outdoor scenes.
+Figure 9 shows two examples from Tanks and Temples Dataset [57] and MultiScan [58] dataset.
+Since our adopted diffusion model only supports images at $256\times 256$ resolution, we downsample the scene point cloud before reconstruction.
+Also, our adopted POCO is relatively slow, especially with large-scale input point clouds.
+In the future, we may seek to develop a more large-scale-friendly version of PointDreamer, by adopting higher-resolution diffusion models and geometry reconstruction models designed for big scenes.
+2D inpainting will remain our key idea, with only submodules replaced.
+TABLE V: Geometry evaluation results of our adopted POCO and baseline methods. Note that baseline Texture Field also uses POCO for geometry extraction.
+{NiceTabular}
+c—c—ccc ShapeNet Cat.   Method CD ↓ NC ↑ FS ↑
+Chair   SPR 1.3094 0.9024 0.8710
+NKSR 0.9079 0.8694 0.8188
+POCO (TF & Ours) 0.4349 0.9267 0.9632
+Car   SPR 0.8428 0.8847 0.8945
+NKSR 0.7876 0.8201 0.8128
+POCO (TF & Ours) 0.4407 0.8677 0.9462
+Motirbike   SPR 4.2807 0.7647 0.6587
+NKSR 0.8786 0.6871 0.7249
+POCO (TF & Ours) 0.4423 0.7663 0.9434
+Dataset   Method CD ↓ NC ↑ FS ↑
+GSO   SPR 0.8565 0.9446 0.9265
+NKSR 0.6149 0.9296 0.9006
+POCO (TF & Ours) 0.4424 0.9474 0.9679
+OmniObject3D   SPR 0.6777 0.9511 0.9407
+NKSR 0.6148 0.9424 0.9003
+POCO (TF & Ours) 0.3393 0.9669 0.9866
+TABLE VI: Quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting.
+{NiceTabular}
+c—cccc Geometry Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
+SPR 19.4071 0.8752 0.1777 74.9854
+Depth Inpainting 26.2502 0.9466 0.0717 20.7198
+Our Adopted POCO 26.2565 0.9516 0.0574 4.9326
+TABLE VII: Quantitative results of replacing our inpainting module from DDNM to nearest interpolation, linear interpolation, and DiffPIR.
+{NiceTabular}
+c—cccc Inpainting Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
+Nearest Interpolation 26.1175 0.9457 0.0618 11.0205
+Linear Interpolation 26.1739 0.9474 0.0612 9.4725
+DiffPIR 26.1582 0.9456 0.0652 9.4823
+Our Adopted DDNM 26.2565 0.9516 0.0574 4.9326
+Figure 9: Our reconstruction results on scenes.
+Figure 10: More visual comparisons with baseline methods. Row 1-3: ShapeNetV2 dataset. Row4-6: GSO dataset. Row 7-9: OmniObject3d dataset.
+Figure 11: Visual results with intermediate multi-view images.
+Figure 12: Pipeline of extracting geometry by depth inpainting.
+TABLE VIII: Quantitative results of using different $K$ values (numbers of viewpoints for projection and inpainting) on the motorbike category of ShapeNetCoreV2 dataset.
 More views contribute to a slightly higher reconstruction quality.
@@
 20 On the 20 vertices of a regular icosahedron 21.2013 0.9299 0.0562 28.4213
-Figure 14.
-Visual comparisons of our reconstructed meshes with different $K$ values (numbers of viewpoints for projection and inpainting).
+Figure 13: Visual comparisons of our reconstructed meshes with different $K$ values (numbers of viewpoints for projection and inpainting).
 An insufficient number of views would lead to artifacts in invisible or occluded areas.
-B.2. Visual Comparisons: Impact of Degraded Input Quality
-B.2.1. Anti-Noise Ability Analysis
-Figure 15 presents the visual comparisons of our method and baseline methods’ reconstructed meshes from noisy or clean input point clouds.
-Overall, we have the following observations, which are consistent with the quantitative results in our manuscript:
-  1. (1)
-    As expected, all methods produce higher-quality textured meshes with clean input point clouds compared to noisy ones.
-  2. (2)
-    Our PointDreamer shows a relatively high anti-noise ability, where only a small performance drop is observed when giving noisy input.
-  3. (3)
-    Our PointDreamer with noisy input point clouds shows an even better visual effect compared to baseline methods with clean inputs.
-Figure 15.
-Visual comparisons of our PointDreamer’s and baseline methods’ reconstructed meshes, with noisy or clean point clouds as input.
+Figure 14: Visual comparisons of our PointDreamer’s and baseline methods’ reconstructed meshes, with noisy or clean point clouds as input.
 Our PointDreamer shows a strong anti-noise ability by producing high-quality textures even with noisy input.
-B.2.2. Sparsity Test
-Figure 16 shows the visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input.
-There is a very small performance drop introduced by decreasing the input point number, which can sometimes be hard to notice by human eyes.
-This indicates a relatively high robustness of our method towards varying degrees of input sparsity.
-Figure 16.
-Visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input.
+Figure 15: Visual comparisons of our PointDreamer’s reconstructed meshes with different numbers of points as input.
 There is a small performance drop when adopting sparser input, which sometimes can be hard to notice by human eyes.
-B.3. Quantitative Results of Method Component Analysis
-B.3.1. Geometry Extraction
-We provide the quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting in Table B.3.1.
-As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID.
-This is consistent with the visual comparisons (see Figure 9) in our manuscript where both SPR and Depth Inpainting suffer from noisy geometry and thus low-quality textures.
-Table 7. Quantitative results of replacing our geometry extraction module from POCO to SPR and Depth Inpainting.
-{NiceTabular}
-c—cccc Geometry Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
-SPR 19.4071 0.8752 0.1777 74.9854
-Depth Inpainting 26.2502 0.9466 0.0717 20.7198
-Our Adopted POCO 26.2565 0.9516 0.0574 4.9326
-B.3.2. 2D Inpainting
-We provide the quantitative results of replacing our 2D inpainting module from DDNM (Wang et al., 2022b) to nearest interpolation, linear interpolation, and DiffPIR (Zhu et al., 2023) in Table B.3.2.
-DiffPIR is another 2D-diffusion-based image restoration approach with inpainting ability.
-As can be seen, DDNM outperforms the other three methods thanks to its strong inpainting ability, which is consistent with the visual comparisons (see Figure 10) in our manuscript.
-Table 8. Quantitative results of replacing our inpainting module from DDNM to nearest interpolation, linear interpolation, and DiffPIR.
-{NiceTabular}
-c—cccc Inpainting Module PSNR ↑ SSIM ↑ LPIPS ↓ FID ↓
-Nearest Interpolation 26.1175 0.9457 0.0618 11.0205
-Linear Interpolation 26.1739 0.9474 0.0612 9.4725
-DiffPIR 26.1582 0.9456 0.0652 9.4823
-Our Adopted DDNM 26.2565 0.9516 0.0574 4.9326
+Figure 16: Visual comparisons of the Texture Field baseline trained with different losses.
+“2D per-pixel” denotes rendering the generated textured mesh to multi-view 2D images, and then calculating the MSE loss between the rendered and GT images.
+“3D per-point” denotes calculating MSE loss between the predicted and GT colors of sampled 3D points.

### v3

@@
 Abstract
-Reconstructing textured meshes from colored point clouds is an important but challenging task.
-Most existing methods yield blurry-looking textures or rely on 3D training data that are hard to acquire.
-Regarding this, we propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud via diffusion-based 2D inpainting.
-Specifically, we first reconstruct an untextured mesh.
-Next, we project the input point cloud into 2D space to generate sparse multi-view images, and then inpaint empty pixels utilizing a pre-trained 2D diffusion model.
-After that, we unproject the colors of the inpainted dense images onto the untextured mesh, thus obtaining the final textured mesh.
-This project-inpaint-unproject pipeline bridges the gap between 3D point clouds and 2D diffusion models for the first time.
-Thanks to the powerful 2D diffusion model pre-trained on extensive 2D data, PointDreamer reconstructs clear, high-quality textures with high robustness to sparse or noisy input.
-Also, it’s zero-shot requiring no extra training.
-In addition, we design Non-Border-First unprojection strategy to address the border-area inconsistency issue, which is less explored but commonly-occurred in methods that generate 3D textures from multiview images.
-Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets show the SoTA performance of PointDreamer, by significantly outperforming baseline methods with 30% improvement in LPIPS score (from 0.118 to 0.068).
+Faithfully reconstructing textured meshes is crucial for many applications.
+Compared to text or image modalities, leveraging 3D colored point clouds as input (colored-PC-to-mesh) offers inherent advantages in comprehensively and precisely replicating the target object’s 360° characteristics.
+While most existing colored-PC-to-mesh methods suffer from blurry textures or require hard-to-acquire 3D training data, we propose PointDreamer, a novel framework that harnesses 2D diffusion prior for superior texture quality.
+Crucially, unlike prior 2D-diffusion-for-3D works driven by text or image inputs, PointDreamer successfully adapts 2D diffusion models to 3D point cloud data by a novel project-inpaint-unproject pipeline.
+Specifically, it first projects the point cloud into sparse 2D images and then performs diffusion-based inpainting.
+After that, diverging from most existing 3D reconstruction or generation approaches that predict texture in 3D/UV space thus often yielding blurry texture, PointDreamer achieves high-quality texture by directly unprojecting the inpainted 2D images to the 3D mesh.
+Furthermore, we identify for the first time a typical kind of unprojection artifact appearing in occlusion borders, which is common in other multiview-image-to-3D pipelines but less-explored.
+To address this, we propose a novel solution named the Non-Border-First (NBF) unprojection strategy.
+Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets demonstrate that PointDreamer, though zero-shot, exhibits SoTA performance ( 30% improvement on LPIPS score from 0.118 to 0.068), and is robust to noisy, sparse, or even incomplete input data.
 Code at: https://github.com/YuQiao0303/PointDreamer.
@@
 1 Introduction
-3D point clouds and meshes are two of the most fundamental 3D modalities in computer vision, graphics, and robotics.
-Reconstructing 3D textured meshes from colored point clouds (abbr.
+Faithful 3D mesh reconstruction is a fundamental task in computer vision and graphics.
+It aims to produce accurate digital recreation of the full geometry and appearance of real-world objects.
+While 3D generation [1, 2, 3] or mesh texturing models [4, 5, 6] from text or image inputs have seen remarkable advances, they produce results that exhibit only coarse semantic alignment or viewpoint-specific consistency compared to the target object.
+As a result, they cannot achieve 360° physically accurate reconstructions with high fidelity.
+In contrast, 3D colored point clouds provide direct measurements of an object’s surface geometry and colors from all perspectives.
+This essential feature theoretically provides a unique foundation for faithful reconstruction.
+Consequently, reconstructing coherent and visually appealing 3D textured meshes from sparse and unstructured colored point clouds (abbr.
 colored-PC-to-mesh) is a pivotal task, with extensive applications like digital twin [7], metaverse [8], cultural heritage preservation [9], etc.
-It requires the accurate reconstruction of both geometry and colors, to transform sparse and unstructured data into coherent and visually appealing models.
-While recent methods [4, 5, 6] have achieved relatively satisfactory geometry reconstruction quality, color reconstruction remains challenging, especially for low-density point clouds.
+Different from 3D generation based on text or images, point cloud reconstruction faces its own challenges.
+While recent point cloud reconstruction methods [10, 11, 12] have achieved relatively satisfactory geometry reconstruction quality, color reconstruction remains challenging, especially for low-density point clouds.
 Unlike images that provide dense pixel-level representations, the inherent sparse nature of point clouds makes it difficult to reconstruct dense colors.
@@
 This inspires us to adopt 2D diffusion models for the colored-PC-to-mesh task, to address the low-quality color reconstruction issue.
-However, it is not trivial to implement this idea: (1) Existing 2D-diffusion-for-3D methods typically focus on 3D generation conditioned on text [15, 16] or images [17, 18], leaving point-cloud-based reconstruction unexplored.
-(2) Existing 2D-diffusion-for-3D methods often yield blurry appearances [19] by predicting colors in 3D space as a texture [20, 16]/radiance [21, 22, 23] field or 3D Gaussians [19, 18].
+However, it is not trivial to implement this idea: (1) Existing 2D diffusion models [20, 21] are predominantly developed to be conditioned on text or image inputs, fundamentally incompatible with point cloud data.
+Consequently, existing 2D-diffusion-based 3D generation methods are driven typically by text [22, 23] or images [3, 24], leaving point-cloud-based reconstruction under-explored.
+(2) Existing 2D-diffusion-for-3D methods often yield blurry appearances [25] by predicting colors in 3D space as a texture field [26, 23], a radiance field [27, 28, 29] or 3D Gaussians [25, 24].
 How to leverage 2D reference images to generate high-quality textures requires more attention.
@@
 Finally, we directly unproject the 2D images onto the mesh.
-Unlike most existing methods that predict colors in 3D [10, 11] or UV spaces [8], directly adopting colors of 2D diffusion models’ output high-quality 2D images leads to clear textures.
+Unlike most existing methods that predict colors in 3D [16, 17] or UV spaces [14], directly adopting the colors of 2D diffusion models’ output high-quality 2D images leads to clear textures.
 In particular, our designed NBF unprojection strategy effectively avoids artifacts caused by inconsistencies in occlusion border regions.
 Overall, we list our contributions below:
-  • We propose PointDreamer, a zero-shot SoTA framework that fills the gap in the less-explored but important task of 3D textured mesh reconstruction from colored point cloud, pioneering in utilizing 2D diffusion models for 3D point cloud reconstruction.
-  • PointDreamer reconstructs high-quality and clear-looking textures for meshes, avoiding the common blurring issue.
-PointDreamer is also zero-shot.
-While requiring no extra training, it shows high generalization capability: robust to noisy, sparse, or even incomplete input points, benefiting from the power of 2D diffusion models.
-  • We design Non-Border-First (NBF) Unprojection, a novel unprojection strategy to apply colors of 2D multiview images onto 3D meshes.
-It effectively avoids artifacts caused by inconsistent prediction near occlusion borders in any multiview-images-to-3D methods.
+  • We propose PointDreamer, a SoTA framework for 3D textured mesh reconstruction from colored point cloud, with multiple advantages: high-quality texture, zero-shot, and robust to noisy, sparse, or even incomplete input data.
+  • Existing 2D diffusion models are predominantly developed for text or image inputs, thus confining existing 2D-diffusion-for-3D methods to these inputs too.
+Unlike them, PointDreamer resolves the intrinsic incompatibility between 2D diffusion and colored 3D point clouds, by a novel project-inpaint-unproject pipeline.
+  • Many existing 3D reconstruction or generation methods yield blurry texture by predicting colors in 3D or UV space.
+In contrast, PointDreamer achieves high-quality texture by directly unprojecting predicted 2D images to 3D.
+  • We identify for the first time a common but less-explored artifact in multiview-image-to-3D pipelines caused by inconsistent prediction near occlusion borders.
+We propose a novel strategy named Non-Border-First (NBF) unprojection to effectively address this issue.
 Experiments on various benchmarks show the SoTA performance of PointDreamer, by significantly outperforming baseline methods quantitatively and qualitatively.
@@
 This may limit its performance and generalization capability, compared to our adopted zero-shot 2D-diffusion-based inpainting approach trained with extensive 2D data.
-2.2 2D Diffusion Models for 3D Generation
-To capitalize on the abundant availability of 2D data and the high performance of 2D diffusion models, researchers seek to “lift” results of 2D diffusion models to 3D.
-Most of these works optimize or learn a NeRF [22, 23], a DMTet with texture field [29, 16], or 3D Gaussianss by Score Distillation Sampling [13] or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model.
-View-conditioned diffusion models [22] or multi-view diffusion models [30, 14, 31] are developed by fine-tuning existing diffusion models for generating multi-view-consistent images.
-Most of these works focus on 3D generation from text or images instead of point clouds.
-They tend to generate blurry textures, and may suffer from artifacts in occlusion border areas.
+2.2 2D Diffusion for 3D Generation and Texturing
+With the success of 2D diffusion models, researchers have explored utilizing them for 3D generation and 3D mesh texturing.
+Most of these generation works optimize or learn a NeRF [28, 29], a DMTet with texture field [2, 35, 23], or 3D Gaussianss by Score Distillation Sampling [1] or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model.
+View-conditioned diffusion models [28] or multi-view diffusion models [36, 19, 37] are developed to produce multiview-consistent images for subsequent 3D generation or texturing.
+Later, Large Reconstruction Models [38, 39, 3] are proposed, demonstrating the effectiveness of directly using feed-forward networks for 3D generation.
+In addition, inpainting-based methods are popular in the field of generating texture for 3D meshes [4, 5, 6].
+They use geometry-aware diffusion models to progressively inpaint the complete texture.
+However, conditioned on only text or a single image instead of a point cloud, the above methods align more closely with “generation” instead of “reconstruction”.
+In other words, they theoretically cannot faithfully reconstruct the target object; see Figure 2 for an example.
+Also, they may generate blurry textures as shown in “CRM” in Figure 2, and may suffer from artifacts in occlusion border areas.
+Figure 2: Comparison with different 2D-diffusion-for-3D methods.
 2.3 Border-Area-Inconsistency Issue in 3D Texturing
@@
 Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training.
-It takes only about 61s and 30s for PointDreamer’s DDNM and nearest/linear interpolation variants (see Figure 7 (b)) to reconstruct a shape of 30k points on an NVIDIA A100 GPU, compared to other 2D-diffusion-for-3D methods such as Zero-1-to-3 [22] (~20 m), DreamGaussin [19] (~2 m), Wonder3D [32] (2~3 m), LGM [18] (~65 s), etc.
-Figure 2: Illustration of hidden point removal, without which the results would be messed up.
-Figure 3: Top: Illustration of input and output of the unprojection module, which unprojects colors of 2D multi-view images to a 3D mesh.
-Middle Left: An naive unprojection implementation.
-For each 3D point $\mathbf{p}$, from all views where $\mathbf{p}$ is visible, the highest-direction-priority view is chosen to paint $\mathbf{p}$.
-Example point $\mathbf{p}$’s corresponding pixel in the selected view $a$ is with incorrect color since it’s near an occlusion border, producing an artifact.
-Bottom Right: our proposed Non-Border-First Unprojection.
-Observing the correspondence between the occlusion border in 2D image space and the visibility border in UV space (both marked by sky-blue dashed lines), we prioritize non-border areas in UV space to avoid such artifacts.
+It takes only about 61s for PointDreamer to reconstruct a shape of 30k points on an NVIDIA A100 GPU, compared to other 2D-diffusion-for-3D methods such as Zero-1-to-3 [28] (~20 m), DreamGaussin [25] (~2 m), Wonder3D [40] (2~3 m), LGM [24] (~65 s), Texture [4] (~5 m), Text2Tex [5] (~15 m), Easi-Tex [6] (~6 m), and ProlificDreamer [2] (several hours).
+Figure 3: Hidden point removal aims to avoid significant stippled artifacts
+Figure 4: The unprojection module takes an untextured mesh and corresponding multiview images as input, to output the associated textured mesh.
+Figure 5: The naive unprojection strategy chooses one best view for each texture atlas pixel by considering only the direction priority (similarity between the 3D point normal direction and the camera direction).
+This can lead to border-area artifacts.
+The optimization-based unprojection strategy uses multiview images as a reference to optimize the texture atlas with per-pixel MSE loss.
+However, the artifacts still exist.
+Figure 6: Illustration of the proposed NBF unprojection strategy.
+By painting each view’s visible texture atlas, we can detect its visibility border areas, which correspond to occlusion borders in the corresponding view’s 2D image.
+Then paint the texture by prioritizing non-border areas.
 3.1 Geometry Extraction: Point to Surface
@@
 Directly projecting the whole point cloud into 2D space would incorporate points from occluded parts that should not be visible from the given view.
-This would mess up the inpainting results and subsequent mesh textures; see Figure 2 for an example.
+This would cause significant stippled artifacts in the inpainting results and subsequent mesh textures; see Figure 3 for an example.
 Therefore, before generating the sparse images, we process the input point cloud by the “Hidden Point Removal” operator [49], which works by transforming the input and extracting the points that reside on its convex hull.
@@
 We compare different $K$ values in our supplementary file.
-Algorithm 1 PyTorch-like pseudocode for Non-Border-First (NBF) strategy
-1# K: number of views
-2# T: texture atlas
-3# I: multiview images
-4# I_dep: [K, I_res, I_res], multi-view depth maps obtained from geometry
-5# T_dep: [K, T_res, T_res], T’s pixels’ corresponding depths in each view
-6# Tuv2Iab: [T_res, T_res, 2]: Mapping T’s pixel coordinate to I’s
-7# chart_mask: [T_res, T_res]: Does T(u,v) belong to a chart (in foreground)?
-8# T_normal: [T_res, T_res, 3]: T’s pixels’ corresponding 3D point’ normal directions
-9# view_dir: [K, 3]: camera view directions
-10# I_rgb: [K, I_res, I_res, 3], multi-view RGB images
-11T = torch.zeros(T_res, T_res, 3)
-12vis_T = calc_vis(I_dep, T_dep, Tuv2Iab) # [K, T_res, T_res], bool
-13vis_T_edges = detect_edge(vis_T) # [K, T_res, T_res], bool
-14chart_edges = detect_edge(chart_mask) # [T_res, T_res], bool
-15border_edges = vis_T_edges - chart_edges.unsqueeze(0).repeat(K)
-16border_areas = dilate(border_edges, kernel) # [K, T_res, T_res], bool
-17cand_view_T_map = non_border_areas = ~border_areas # [K, T_res, T_res]
-18Tuv_cand_view_num = non_border_areas.sum(0) # [T_res, T_res], int
-19cand_view_T_map[:, Tuv_cand_view_num==0] = vis_T[:, Tuv_cand_view_num==0]
-20dir_priority = calc_priority(T_normal, view_dir) # [K, T_res, T_res]
-21dir_priority[~cand_view_T_map] -= 1000 # Non-Border-First
-22uv_view_id = dir_priority.argmax(0) # [T_res, T_res]
-23for u in range(0, T_res):
-24 for v in range(0, T_res):
-25 T[u, v] = I_rgb[uv_view_id[u, v], Tuv2Iab[u, v, 0], Tuv2Iab[u, v, 1]]
 3.3 Inpainting: Sparse to Dense
@@
 Regarding this, we design a novel approach namely “Non-Border-First Unprojection”.
-The top of Figure 3 illustrates the basic idea of the unprojection module.
+Figure 4 illustrates the basic idea of the unprojection module.
 Specifically, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$.
 We follow existing works [50, 40] to conduct UV mapping by Xatlas [51].
-A texture atlas is an RGB image containing different charts.
+A texture atlas is an RGB image containing different charts (e.g.
+①, ② … in Figure 4).
 Each chart is a continuous area in UV space that corresponds to a continuous segment of $\mathcal{M}$ in 3D space.
+For example, in Figure 4, chart ① corresponds to the bottom side of the brown seat, and chart ② corresponds to the bottom side of the gray cone.
 Each pixel $\mathcal{T}[u,v]$ of a chart corresponds to a surface point $\mathbf{p}$ of $\mathcal{M}$.
@@
 3.4.1 Naive Unprojection: Only Use Direction Priority
-An intuitive idea to pick the best view for $\mathbf{p}$ or $\mathcal{T}[u,v]$ is by direction priority, as shown in the middle left part of Figure 3.
+An intuitive idea to pick the best view for $\mathbf{p}$ or $\mathcal{T}[u,v]$ is by direction priority, as shown in ‘Naive Unprojection Strategy’ on top of Figure 5.
 Specifically, for each view $k$ where point $\mathbf{p}$ can be seen, we calculate the direction priority score $s_{\mathbf{p}}(k)$, defined as the cosine similarity of $\mathbf{p}$’s normal direction $\mathbf{n}$ and view $k$’s camera direction $\mathbf{d}_{k}$.
 The view with the highest direction priority is chosen to paint $\mathbf{p}$ and $\mathcal{T}[u,v]$.
-For example, in the middle left part of Figure 3, $\mathbf{p}$ is painted to gray according to $I_{a}[i_{a},j_{a}]$ since view $a$ has a higher direction priority.
+For example, in the top row of Figure 5, $\mathbf{p}$ is painted as gray according to $I_{a}[i_{a},j_{a}]$ since view $a$ has a higher direction priority.
 Also, if there are a few points invisible from any view, we can still assign a view by direction priority.
-While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the middle left part of Figure 3, where a gray artifact is visible in the reconstructed textured mesh that should have been brown.
-Through analysis, such artifacts appear in border areas: the area near the border of occluded and occluding part of the shape when viewing from view $k$.
+While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the top row of Figure 5, where a gray artifact is visible in the reconstructed textured mesh that should have been brown.
+As shown in ‘Analysis’ in Figure 5, such artifacts appear in border areas: the area near the border of the occluded and occluding part of the shape when viewing from view $k$.
 For example, in view $a$, the brown seat’s bottom side is occluded by the gray bottom of the chair.
@@
 In such areas, the inpainting module and the geometry extraction module may inconsistently delineate the boundary separating the two parts.
-For example, the gray bottom is larger in the inpainted image $I_{a}$ than of the reconstructed mesh, thus resulting in an artifact.
+For example, the gray bottom is larger in the inpainted image $I_{a}$ than of the reconstructed mesh, thus resulting in the artifact.
 3.4.2 Optimization-based Unprojection
 Instead of selecting one single view by direction priority, 3D generation method DreamGaussian [25] adopts an optimization-based strategy to get the textured mesh from multi-view images.
-The idea is to optimize the texture atlas by a per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh.
-In this way, the optimization process kinds of fuses corresponding colors in different views instead of selecting the best one.
+As shown in ‘Optimization Unprojection Strategy’ of Figure 5, the idea is to optimize the texture atlas by a per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh.
+In this way, the optimization process fuses corresponding colors in different views instead of selecting the best one.
 However, it cannot ignore the inconsistent colors in the border areas, so the artifacts can only be reduced instead of eliminated.
@@
 3.4.3 Non-Border-First Unprojection (NBF)
-The bottom of Figure 3 illustrates the idea of our proposed novel unprojection strategy namely Non-Border-First Unprojection.
-It begins with detecting border areas in multi-view images, so as to paint the texture atlas while deprioritizing them.
-After that, if some atlas pixels remain unpainted since their associated 3D points can not be seen from any non-border area, we then select from all areas, whether they are border areas or not, to paint the rest pixels.
-Now the question is, how to accurately detect border areas.
-Interestingly, we find that, the borders between occluded and occluding parts of the shape in image $I_{k}$ (2D space) often correspond to the borders between visible and invisible areas in “view $k$ visible texture atlas $\mathcal{T}_{k}$” (UV space), where $\mathcal{T}_{k}$ is defined as a texture atlas in which only visible areas in view $k$ are filled and rest invisible areas are empty.
-See the bottom of Figure 3 as an example, where such border areas are marked by sky blue dashed lines.
-In posed images $I_{a}$ and $I_{b}$, these areas divide the occluded brown seat and the occluding gray bottom, while in view $a$ and $b$ visible texture atlas $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$, these areas divide the visible (painted) and invisible (transparent) parts in view $a$ and $b$, respectively.
-The reason for this correspondence of UV visibility border and 2D occlusion border is that, a continuous area within a chart of the texture atlas usually corresponds to a continuous segment of the 3D surface.
+Building upon the above analysis, we propose the Non-Border-First (NBF) Unprojection strategy to eliminate artifacts arising from inconsistent predictions in border areas.
+The core principle is to prioritize assigning colors from non-border regions across views when filling the texture atlas.
+Border Area Detection.
+To prioritize non-border areas, we first need to detect border areas in each view.
+Since the direct detection of occlusion borders within multiview 2D images is challenging, we propose to shift the detection process to UV space.
+This approach leverages a key correspondence we observe empirically: occlusion borders in image $I_{k}$ (2D space) often correspond to visibility borders in “view $k$ visible texture atlas $\mathcal{T}_{k}$” (UV space), where $\mathcal{T}_{k}$ is defined as a texture atlas in which only visible areas in view $k$ are filled and rest invisible areas are empty.
+For example, Figure 6 (1) illustrates view $a$ and view $b$’s visible texture atlas $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$, respectively.
+As the navy blue dashed arrows show, the bottom side of the brown seat in $I_{a}$ and $I_{b}$ are mainly used to paint corresponding pixels in chart ① of $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$, and the bottom side of the gray cone is mainly used to paint chart ②.
+Zooming in, we can see in Figure 6 (2) that, the representive chart ① in both $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ can be divided into a brown area (visible) and an empty area (invisible), respectively.
+We mark their border areas in sky blue dashed lines and denote such borders as “visibility border”.
+Note that the brown area corresponds to the bottom side of the brown seat of the chair, and the empty area corresponds to the parts that are occluded by the gray cone, as we marked by arrows with text “Correspond!” in sky blue in Figure 6.
+In other words, the invisible/visible areas in $\mathcal{T}_{k}$ correspond to the occluded/occluding areas in $I_{k}$.
+The reason for this correspondence is that, a continuous area within a chart of the texture atlas usually corresponds to a continuous segment of the 3D surface.
 When a chart contains both visible and invisible regions in a given view $k$, it often indicates that the corresponding 3D areas of invisible regions are occluded by some other part of the shape.
-Therefore, we design our Non-Border-First unprojection strategy as follows.
+Border-Priority-Based Unprojection.
+Exploiting this correspondence, visibility borders in $\mathcal{T}_{k}$ (detected via edge detection algorithms) serve as proxies for occlusion borders in $I_{k}$.
+Consequently, NBF prioritizes non-border regions during texturing.
+For instance, in Figure 6 (3), example 3D point $\mathbf{p}$ and its corresponding pixel $I_{a}[i_{a},j_{a}]$ and $I_{b}[i_{b},j_{b}]$ in $I_{a}$ and $I_{b}$ are marked in small red circles.
+Since $I_{a}[i_{a},j_{a}]$ and $I_{b}[i_{b},j_{b}]$ are in border and non-border areas, respectively, the non-border-first principle guides us to pritize $I_{b}[i_{b},j_{b}]$ and paint point $\mathbf{p}$ as $I_{b}[i_{b},j_{b}]$’s color (brown).
+In this way, we effectively eliminate the gray artifact in naive or optimization unprojection caused by inconsistent prediction in border areas.
+Putting everything together, the detailed steps of the proposed NBF unprojection strategy are as follows.
   1. 1.
@@
   5. 5.
-    If some points cannot be seen from either view, we assign it with a view by direction priority considering all areas in all views whether they’re border areas or not.
+    If some points cannot be seen from either view, we assign each of them with a view by direction priority considering all areas in all views whether they’re border areas or not.
   6. 6.
     We experimentally find that additionally optimizing from the generated texture atlas by only calculating the loss of non-border pixels would further enhance the results.
-In this way, our non-border-first strategy can reasonably eliminate artifacts caused by inconsistent prediction in border areas.
-Altorithm 1 shows the corresponding pseudoCode.
-We conduct experiments to verify the effectiveness of our proposed strategy in Section 4.
 4 Experiments and Results
@@
 Since we focus on texture instead of geometry reconstruction, we put geometry evaluation results in our supplementary file.
-4.2 Baselines
+4.1.3 Baselines
 To fully evaluate the performance of our method, we compare it with three kinds of baseline methods:
@@
 4.2 Main Results
-4.3.1 Qualitative Results
+4.2.1 Qualitative comparisons
 We present visual results in Figure 7.
-Since DHSP3D is much slower than other methods (4-6 hours per shape compared to $<$1 min), we randomly select two objects from GSO dataset for efficiency during comparison.
+Since DHSP3D is much slower than other methods (4-6 hours per shape compared to <<1 min), we randomly select two objects from GSO dataset for efficiency during comparison.
 We highly recommend referring to our supplementary materials for more visual results and a 360° video.
@@
 4.2.2 Quantitative comparisons
-We summarize the quantitative results in Tables 4.2, and geometry evaluation results in our supplementary file.
+We summarize the quantitative results in Tables 4.1.3, and geometry evaluation results in our supplementary file.
 Our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin.
@@
 While achieving visually better results, a few pixels with extremely high errors may lower the PSNR score.
-4.4 Multi-view Consistency
+4.2.3 Multi-view Consistency
 Figure 7 (e) and our supplementary video show that, our method achieves high multi-view consistency.
 The reason is two-fold: (1) Unlike image-conditioned 3D generation, it’s easier to maintain multi-view consistency when reconstructing point clouds, since the input point cloud already determines the global shape and colors, leaving only local details to be dreamed by the model.
-(2) For inconsistencies in local details, our proposed Non-Border-First unprojection effectively addresses this issue; see Figure 7 (c) and Table 4.4.
+(2) For inconsistencies in local details, our proposed Non-Border-First unprojection effectively addresses this issue; see Figure 11 (c) and Table 4.2.4.
+4.2.4 More Visual Presentation
+For readers to better understand our pipeline, we present the input point cloud, intermediate multiview images, and output meshes in Figure 8.
+Figure 8: Visual presentation of the input point clouds, intermediate sparse and dense multiview images, and the output meshes.
 TABLE II: Anti-noise comparisons on ShapeNetCoreV2 Chairs. ‘Noisy’ means adding Gaussian noise with a standard deviation of 0.005 to the input.
@@
 This is benefited from the strong inpainting power of the adopted 2D diffusion model, as illustrated in (b).
-Figure 7: Sub-Module Replacement Results: (a) Geometry module. (b) Inpainting module. (c) Unprojection module.
+Figure 11: Ablation comparisons between our proposed NBF strategy and other unprojection strategies.
 Figure 12: Examples of adopting our proposed NBF strategy on other methods (InstantMesh [59]).
@@
 We compare our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset.
-Table 4.4 shows the results, where our method, even with noisy input, outperforms baselines with clean input.
+Table 4.2.4 shows the results, where our method, even with noisy input, outperforms baselines with clean input.
 Figure 9 (b) also shows our PointDreamer’s strong anti-noise ability.
-While baseline methods exhibit significant performance drops given noisy input, our PointDreamer’s performance degradation is minimal: even barely noticeable by human perception.
+While baseline methods exhibit significant performance drops given noisy input, our PointDreamer’s performance degradation is minimal, even barely noticeable by human perception.
+We analyze that PointDreamer’s relative robustness to noisy input stems from the strong diffusion prior.
+Although the input sparse images are noisy, the adopted 2D diffusion model is trained on a vast amount of clean and plausible 2D images with no noise.
+Consequently, the model yields relatively clean inpainted images as it learns to generate samples that follow its training data’s distribution.
 4.3.2 Sparsity Test
-To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 10k gradually, and present the results in Table 4.4.
+To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 10k gradually, and present the results in Table 4.2.4.
 We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input.
@@
 We also provide visual comparisons in Figure 9 (c), where our method significantly outperforms baseline methods with only 10k points as input.
+We analyze that the high generation capability of the 2D diffusion prior is the main reason for this robustness to sparse input data.
 4.3.3 Texture Completion Ability
@@
 As a result, as shown in Figure 10 (c), our PointDreamer completes the missing area with high-quality texture compared to other methods.
-4.6 Method Component Analysis
-We here compare different implementations for sub-modules (geometry extraction, inpainting, and unprojection module) of our method by replacing one module at a time while keeping others unchanged.
-All experiments in this subsection are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
-4.6.1 Geometry Extraction
-In addition to POCO, we further explore extracting geometry by SPR and depth inpainting to investigate how geometry quality influences the final reconstructed textured mesh.
-Please refer to our supplementary file for more details on depth inpainting and quantitative results.
-Figure 7 (a) shows that noisy geometry leads to noisy texture, and our adopted POCO performs the best.
-4.6.2 2D Inpainting
-We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR [53], and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability.
-We provide the visual comparisons in Figure 7 (b) and quantitative results in our supplementary file.
-As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, indicating the importance of a strong inpainting module to the reconstruction performance.
-4.6.3 Unprojection
-We present the comparison results of different unprojection strategies in Figure  7 (c) and Table 4.4.
-  • Opt. Scratch: optimize the texture atlas from scratch;
+4.4 Ablation Study on NBF Unprojection
+To verify the effectiveness of our proposed NBF unprojection strategy, we conduct comparison experiments on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
+We present the comparison results in Figure 11 and Table 4.2.4, including the following methods:
+  • Opt.
+Scratch: Randomly initialize a texture atlas and then optimize it by minimizing the per-pixel MSE loss between the mesh renderings and the inpainted multiview images, as described in Section 3.4.2;
   • Naive: select the best views by direction priority;
   • NBF: select the best views by our proposed Non-Border-First strategy;
-  • Opt. Naive: further optimize Naive’s obtained atlas;
-  • Opt. NBF: further optimizes NBF’s obtained atlas while considering only loss in non-border areas.
-Figure 7 (c) shows that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization.
+  • Opt.
+Naive: Take Naive’s obtained atlas as initialization and optimize it by minimizing the per-pixel MSE loss between the mesh renderings and the inpainted multiview images, as described in Section 3.4.2;
+  • Opt. NBF: Take NBF’s obtained atlas as initialization and optimize it by minimizing the per-pixel MSE loss between the mesh renderings and the inpainted multiview images’ non-border areas.
+Figure 11 shows that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization.
 Only our NBF and Opt.
 NBF nearly eliminate such artifacts.
-Table 4.4 also shows that our proposed NBF and Opt.
+Table 4.2.4 also shows that our proposed NBF and Opt.
 NBF outperform other unprojection strategies.
-Our NBF strategy can also work for other methods [51, 37] that reconstruct mesh textures from multiview images.
-For example, InstantMesh [51] reconstructs a textured 3D mesh from a single image, by first generating multiview images, and then predicting the mesh by a large reconstruction model.
-As mentioned, predicting colors by a network trained with 3D data can result in blurry textures, as shown in Figure 8 (a).
-While unprojecting the multiview images to the mesh can achieve clearer textures, as shown in Figure 8 (b), artifacts often occur at border areas, such as the border of the eye as marked in red.
-Our NBF strategy effectively solved such artifacts, as shown in (c).
+Importantly, going beyond the specific colored-PC-to-mesh task, our proposed NBF unprojection strategy can effectively generalize to any method that textures an existing mesh given multiview images.
+In these methods, even with specifically designed techniques, the local border-area inconsistencies between multiview images and geometry can hardly be eliminated, which our NBF addresses.
+We take InstantMesh [59] as an example.
+It reconstructs a textured 3D mesh from a single image, by first generating multiview images through a multiview diffusion model [61], and then outputting the mesh via a large reconstruction model that directly predict the mesh’s geometry and texture in 3D space.
+Note that the multiview diffusion technique is specifically designed for multiview-consistent generation.
+We present visual comparisons in Figure 12.
+As mentioned in our introduction, learning colors in 3D space can result in blurry textures, as shown in Figure 12 (a).
+When unprojecting the multiview images to the mesh using the naive unprojection strategy, the texture becomes sharper, as shown in Figure 12 (b).
+However, as a cost, an artifact occurs in a border area, i.e.
+the border of the eye as marked in red.
+Our NBF strategy effectively solves this artifact, as shown in (c).
 More examples are provided in (d-g).
+These examples demonstrate that NBF can serve as a general solution to border-area artifacts in multiview-image-to-texture methods.
+4.5 Real-life Data Experiments
+Above, we followed the common routine of point cloud reconstruction works [47, 14, 10] to conduct experiments using point clouds sampled from existing meshes, because the calculation of existing quantitative evaluation metrics like PSNR require ground-truth meshes.
+To better evaluate the effectiveness of our PointDreamer, we further capture point cloud inputs by scanning real-life objects with an Intel RealSense L515 LiDAR sensor.
+To ensure a more comprehensive evaluation, we select objects with diverse shapes and scales.
+As shown in Figure 13, the bottle in the top row is approximately a cylinder, the “Aoyu” in the second row and the box in the bottom row mainly consist of planes, and the bag in the third row exhibits a less regular shape.
+The bottle is the smallest in real life, leading to the noisiest input point cloud due to the limited sensor resolution.
+Since no ground-truth meshes are available for quantitative assessment of these scanned objects, we present the qualitative results in Figure 13.
+Figure 13: Comparison results with real-scanned point clouds as input. Better zoom in.
+As shown in Figure 13, other methods suffer from different levels of blurriness or jagged effects.
+Compared to them, our PointDreamer archives clearer texture reconstruction, thanks to the diffusion prior.
 5 Conclusion, Limitations and Future Work
@@
 Appendix A Geometry Evaluation Results
-To compare the reconstructed geometry quality of PointDreamer and baseline methods while ignoring textures, we report geometry evaluation results in Table F.
+To compare the reconstructed geometry quality of PointDreamer and baseline methods while ignoring textures, we report geometry evaluation results in Table H.
 Same as our manuscript, we use ShapeNetCoreV2 [52] (Chair, Car, Motorbike categories), Google Scanned Objects [53] and Omniobject3d [54] datasets.
@@
 Figure 16 shows more visual comparisons of our PointDreamer against baseline methods.
-Figure 11 shows more visual results of our PointDreamer, including projected sparse multi-view images and inpainted dense multi-view images.
-We hope this can help the readers better understand our pipeline.
 Appendix C Method Component Analysis
+We here compare different implementations for our geometry extraction and inpainting sub-modules by replacing one module at a time while keeping everything else unchanged.
+All experiments in this section are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
 C.1 Geometry Extraction Module Replacement Experiment
@@
 C.1.2 Results
-We present the visual comparisons in our manuscript and quantitative results in Table F, respectively.
+We present the visual comparisons in Figure 14 (a) and quantitative results in Table H, respectively.
 As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID.
@@
 C.2 2D Inpainting Module Replacement Experiment
-We provide the quantitative results of replacing our 2D inpainting module from DDNM [24] to nearest interpolation, linear interpolation, and DiffPIR [53] in Table F.
-DiffPIR is another 2D-diffusion-based image restoration approach with inpainting ability.
-As can be seen, DDNM outperforms the other three methods thanks to its strong inpainting ability, which is consistent with the visual comparisons from Figure 4 (b) in our manuscript.
+We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR [65], and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability.
+We provide the visual comparisons in Figure 14 (b) and quantitative results in Table H.
+As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, leading to a lower quantitative score.
+This indicates the importance of a strong inpainting module to the reconstruction performance.
 Appendix D More Experimental Results
@@
 To investigate the effect of different numbers of viewpoints (denoted as $K$) for projection and inpainting, we conduct experiments on the motorbike category or ShapeNetCoreV2 dataset by using different $K$ values including 6, 8, and 20.
-Table F shows the distribution of cameras for each setting, together with the quantitative results.
+Table H shows the distribution of cameras for each setting, together with the quantitative results.
 We can see that more views contribute to a slightly higher reconstruction quality.
@@
 This indicates a relatively high robustness of our method towards varying degrees of input sparsity.
+D.3 Comparison with recent image-to-3D, text-to-3D, and mesh texturing methods
+D.3.1 Experimental Setting
+We select two of the most updated and representative methods from each relevant category (mesh texturing, image-to-3D, and text-to-3D), which results in the following comparison baselines:
+  • StableDreamFusion: A text-to-3D method based on SDS optimization.
+As the original DreamFusion is not open-source, we utilized an unofficial open-source implementation with substantial community support (8.6k GitHub stars).
+  • ProlificDreamer: Another text-to-3D method leveraging SDS and further DMTet optimization.
+  • Text-to-tex: A text-conditioned mesh texturing (text-to-texture) method that employs diffusion-based inpainting.
+  • Easi-tex: An image-conditioned mesh texturing (image-to-texture) method, also utilizing diffusion-based inpainting.
+  • CRM: An image-to-3D method built upon multiview diffusion and a large feed-forward reconstruction network.
+  • StableFast3D (SF3D): Another image-to-3D approach, characterized utilizing large feed-forward reconstruction network.
+The above methods involve three kinds of inputs:
+  • Reference Image: For methods requiring a reference image, we rendered images from the ground-truth mesh of the object.
+  • Untextured Mesh: For mesh texturing methods, we directly utilized the meshes reconstructed by our PointDreamer as the untextured mesh input.
+  • Text Prompt: To ensure fair comparison for text-based methods, we initially generated prompts by feeding the rendered ground-truth images into a large language model (Qwen2.5-VL-32B-Instruct).
+These automatically generated prompts were then manually refined to provide more accurate and detailed descriptions of the objects.
+The complete text prompts are provided below:
+    • Vintage analog clock: dark metal frame, matte khaki face with Arabic numerals.
+Hands at around 10:06:25.
+Small inset dial between center & numeral 12; ‘CROSLEY’ between center & numeral 6.
+Base with two legs, top ring for hanging.
+    • A flat-sided wooden lion board with dual wooden wheels below.
+Features: yellow body, red petal mane, black eyes/orange nose/pink ears/blush marks.
+Green background & handle on lion’s back, ’Fisher-Price’ in white on red at base under the lion.
+Note that most of these methods are relatively time-consuming (from about 15 minutes to more than an hour, compared to our method’s 100 seconds).
+D.3.2 Experimental Results.
+We present the comparison results in Figure 22, which shows:
+  • The text-to-3D methods StableDreamFusion and ProlificDreamer generate significantly different results compared to the ground truth, since they only rely on text for generation without any visual information.
+Even when not considering fidelity, the generation results suffer from severe implausibility (e.g.
+ProlificDreamer’s generated clock) or even missing geometry (e.g.
+StableDreamFusion’s generated clock)
+  • Text-to-tex is free from the geometry unplausibility issue with a mesh provided as input, but still it can hardly faithfully reconstruct the texture.
+  • Easi-tex, though with an input image for reference, still struggles to generate high-fidelity textures, especially the backside without input information.
+  • CRM generates higher-fidelity results then previous methods, but it still can not faithfully reconstruct the back view. In addition, its generated textures are blurry.
+  • SF3D generates clearer textures from the front view, but its generated back views look highly implausible.
+  • Our PointDreamer can reconstruct the objects with the highest fidelity, in other words, most similar to the ground truth object.
+In summary, only our PointDreamer can achieve relatively high fidelity reconstruction, both theoretically and experimentally.
 Appendix E Implementation Details of Texture Field Baseline
@@
 2D inpainting will remain our key idea, with only submodules replaced.
+Appendix G Results on Objaverse Dataset
+We provide some qualitative comparisons on Objaverse dataset in Figure 23.
+Appendix H How UV unwrapping influence NBF
+In the proposed NBF unprojection strategy, a core step is to detect border areas.
+Since we conduct the border detection process in UV space, a natural question is: How does UV unwrapping influence this process?
+For example, what if a continuous region of a 3D mesh is partitioned into two disconnected charts in UV space?
+To answer these questions, we discuss from both experimental and theoretical perspectives:
+  • Our experiments show that Xatlas provides relatively robust UV unwrapping, enabling NBF to outperform baselines both qualitatively and quantitatively, as shown in the experimental results of our manuscript in Section 4.4 “Ablation Study on NBF Unprojection”.
+  • Theoretically, when a single region of a 3D mesh is partitioned into two disconnected charts in UV space, NBF unprojection roughly degenerates to naive unprojection within the additional mis-segmented boundary region.
+Consequently, the lower performance bound of NBF is approximately equivalent to that of the naive unprojection.
+The detailed derivation is provided below.
+When a single region of a 3D mesh is partitioned into two disconnected charts in UV space, this results in an additional segmentation boundary region that now corresponds to “chart borders” (dilated chart edges) in UV space.
+Crucially, for any view where these chart borders are visible, chart borders inherently belong to visibility borders, since they are near empty (invisible) areas that belong to no chart.
+Consequently, the border detection step does not distinguish these views’ priority (now that they are all border areas), and subsequent steps rely solely on their directional priority.
+Therefore, NBF unprojection for such additional segmentation boundary regions becomes equivalent to the naive unprojection without considering border prioritization.
 TABLE V: Geometry evaluation results of our adopted POCO and baseline methods. Note that baseline Texture Field also uses POCO for geometry extraction.
@@
 Our Adopted DDNM 26.2565 0.9516 0.0574 4.9326
+Figure 14: Sub-Module Replacement Results: (a) Geometry module. (b) Inpainting module.
 Figure 15: Our reconstruction results on scenes.
 Figure 16: More visual comparisons with baseline methods. Row 1-3: ShapeNetV2 dataset. Row4-6: GSO dataset. Row 7-9: OmniObject3d dataset.
-Figure 11: Visual results with intermediate multi-view images.
 Figure 17: Pipeline of extracting geometry by depth inpainting.
@@
 “3D per-point” denotes calculating MSE loss between the predicted and GT colors of sampled 3D points.
+Figure 22: Comparison results with recent mesh texturing, image-to-3D, and text-to-3D methods.
+Figure 23: Results on the Objaverse dataset.